### v2

# RG-Invariant Symmetry Ratio for QCD: A Study of $U(1)_{A}$ and Chiral Symmetry Restoration

Abstract
[…]
1 Introduction
The pattern of symmetry realization is a defining feature of any quantum field theory, governing its phase structure and the spectrum of its excitations. In Quantum Chromodynamics (QCD), the theory of the strong interaction, two global symmetries play a pivotal role: the chiral $SU(2)_{L}\times SU(2)_{R}$ symmetry of the light ($u$, $d$) quark sector, and the $U(1)_{A}$ axial symmetry. In the vacuum, the former is spontaneously broken by the chiral condensate $\langle\bar{q}q\rangle$ Nambu:1961tp ; Nambu:1961fr , giving mass to nucleons and generating pions as pseudo-Goldstone bosons. The latter is explicitly broken by the axial anomaly Adler:1969gk ; Bell:1969ts ; Fujikawa:1979ay , contributing significantly to the mass of the $\eta^{\prime}$ meson tHooft:1976snw ; Witten:1979vv ; Veneziano:1979ec . A cornerstone of modern nuclear physics is understanding how these broken symmetries behave under extreme conditions of temperature and density, such as those realized in heavy-ion collisions or within neutron stars.
The restoration of chiral symmetry is associated with the transition from a hadronic phase to a quark-gluon plasma (QGP). For QCD with physical quark masses, this is a smooth crossover occurring at a temperature $T_{c}\sim 156$ MeV Aoki:2006we ; Borsanyi:2013bia ; HotQCD:2014kol ; HotQCD:2018pds ; Borsanyi:2020fev .
A profound and long-standing question is whether the effective restoration of the $U(1)_{A}$ symmetry, linked to the suppression of topological gauge fluctuations, coincides with this chiral crossover or occurs at a distinctly higher temperature $T_{1}>T_{c}$ Pisarski:1983ms ; Leutwyler:1992yt ; Cohen:1996ng ; Aoki:2012yj . Resolving this hierarchy is essential for a complete understanding of the QGP’s structure, the nature of the QCD transition, and the validity of effective models. Progress on this question via first-principles lattice QCD simulations has been challenging. A consistent picture among them has not emerged yet, see e.g., refs. Cossu:2013uua ; Buchoff:2013nra ; Brandt:2016daq ; Tomiya:2016jwr ; Ding:2020xlj ; Kaczmarek:2023bxb ; Aoki:2020noz ; Chiu:2023hnm ; Gavai:2024mcj , and the recent review by H.T. Ding at Lattice 2025222https://indico.global/event/14504/contributions/137953/, 2–8 November, Mumbai, India.
The $U(1)_{A}$ anomaly is particularly sensitive to lattice artifacts, and its clean study requires fermion discretizations that preserve chiral symmetry. Lattice studies with chiral fermions (domain-wall Kaplan:1992bt ; Kaplan:1992sg or overlap Neuberger:1997fp ; Narayanan:1994gw ) have provided crucial insights Cossu:2013uua ; Buchoff:2013nra ; Tomiya:2016jwr ; Aoki:2020noz ; Chiu:2023hnm ; Gavai:2024mcj . Notably, the JLQCD collaboration, using $N_{f}=2$ Möbius domain-wall fermions with reweighting for overlap fermions at the lattice spacing 0.07 fm, found clear evidence that for $T>190$ MeV, the $U(1)_{A}$ breaking is consistent with zero within statistical errors Aoki:2020noz . This is also consistent with studies using $N_{f}=2+1+1$ optimal domain-wall fermions at the physical point and lattice spacing 0.063 fm, where $U(1)_{A}$ axial symmetry is restored for $T\gtrsim 190$ MeV Chiu:2023hnm . However, refs. Aoki:2020noz ; Chiu:2023hnm have not determined the $U(1)_{A}$ symmetry breaking for $T<190$ MeV. On the other hand, in ref. Gavai:2024mcj , using $N_{f}=2+1$ Möbius domain-wall fermions at multiple lattice spacings, the authors observed that the $U(1)_{A}$ axial symmetry is not restored for $T\lesssim 186$ MeV. To investigate whether any discrepancies between these different studies would occur for $T<190$ MeV, a systematic approach to obtain definite continuum-extrapolated results for physical QCD with a chirally symmetric action remains a high-priority goal for the community. Addressing this goal requires overcoming two interconnected challenges: first, performing controlled continuum extrapolations across the temperature range of interest; second, developing a robust, quantitative observable to compare symmetry-breaking strength across different channels.
Traditional probes, such as hadron thermal and screening masses or the behavior of specific correlation functions at a single Euclidean space/time, can be ambiguous or sensitive to analysis choices. What is needed is a renormalization-group (RG) invariant measure that integrates spectral information, provides a clear normalization, and allows for a direct comparison between $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ breaking.
In this paper, we address both challenges. First, we introduce a novel, universal diagnostic: the RG-invariant symmetry ratio $\kappa_{AB}$. For two operators $A$ and $B$ related by a symmetry transformation, we define $\kappa_{AB}=(\chi_{A}-\chi_{B})/(\chi_{A}+\chi_{B})$, where $\chi_{A}\geq\chi_{B}>0$ denote the corresponding susceptibilities derived from the regularized, vacuum-subtracted correlation functions. This construct is bounded $\kappa_{AB}\in[0,1]$, and crucially RG-invariant for exact symmetry partners, making it an ideal model-independent probe.
Second, we present the first application of $\kappa_{AB}$ to the problem of symmetry restoration in QCD. We perform lattice simulations with $N_{f}=2+1+1$ optimal domain-wall quarks at the physical point, using three lattice spacings and twelve temperatures in the range 164–385 MeV. We compute $\kappa_{AB}$ across three distinct symmetry-breaking channels in the nonsinglet sector: the $U(1)_{A}$-breaking (scalar, pseudoscalar) channel ($\kappa_{PS}$) probing the $\pi$-$\delta$ system; the $SU(2)_{L}\times SU(2)_{R}$-breaking (vector, axial-vector) channel ($\kappa_{VA}$) probing the $\rho$-$a_{1}$ system; and an additional $U(1)_{A}$-sensitive channel using tensor–axial-tensor operators ($\kappa_{TX}$) which probes the $\rho_{T}$-$b_{1}$ system through the $\bar{q}\gamma_{4}\gamma_{k}q$ and $\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ currents. The inclusion of the tensor channel provides a vital cross-check, as it probes the $U(1)_{A}$ anomaly through a different Dirac structure than the scalar channel.
Our key findings are as follows. At finite lattice spacing, we observe a clear ordering: the breaking in the $U(1)_{A}$ scalar–pseudoscalar channel is strongest, while the $U(1)_{A}$ tensor–axial-tensor channel and the $SU(2)_{L}\times SU(2)_{R}$ vector–axial-vector channel are significantly weaker and comparable to each other ($\kappa_{PS}>\kappa_{TX}\sim\kappa_{VA}$). However, a controlled continuum extrapolation reveals that this entire hierarchy collapses. All three symmetry-breaking strengths become statistically indistinguishable within our resolution. This result provides a new, high-precision benchmark from a chirally symmetric action. It indicates that the effective restoration scales probed by these different operators converge in the continuum limit, suggesting that the manifestations of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries in the nonsinglet sector are much more concurrent than suggested by finite-lattice-spacing studies focused on a single channel.
The paper is organized as follows. In section 2, we formally define the RG-invariant symmetry ratio $\kappa_{AB}$ and detail its theoretical properties. Our lattice setup, time-correlation functions, and analysis methodology are described in section 3. Numerical results at finite lattice spacing are presented in section 4, and the continuum extrapolation is performed in section 5. The implications of our findings and future applications of the $\kappa_{AB}$ framework are discussed in section 6.
2 RG-invariant symmetry ratio
We introduce a renormalization-group invariant quantity that quantifies the degree of symmetry breaking in a quantum field theory. The construction relies on integrated spectral weights of Euclidean correlation functions for symmetry-related operators. Our measure, the symmetry strength parameter $\kappa_{AB}$, provides a global, scale-free indicator of symmetry violation and will be applied later to chiral symmetry restoration in finite-temperature QCD.
2.1 Correlation functions and symmetry partners
In a theory with an exact symmetry, correlation functions of operators related by the symmetry transformation must be degenerate. Consider the Euclidean time correlation function of a local meson operator
$O_{\Gamma}^{a}(\vec{x},t)=\bar{q}(\vec{x},t)\,\Gamma\,t^{a}\,q(\vec{x},t),$
where $\Gamma$ is a Dirac matrix specifying the quantum numbers of the channel, $t^{a}$ (with $a=1,\dots,N_{f}^{2}-1$) are the generators of $SU(N_{f})$ for flavor nonsinglets, and $t^{0}\equiv\mbox{1\hskip-3.41432ptI}=\text{diag}(1,1,\dots,1)/\sqrt{2N_{f}}$ (identity operator) for the flavor singlet, with the normalization condition ${\rm tr}(t^{a}t^{b})=\delta^{ab}/2$ ($a=0,\dots,N_{f}^{2}-1$). For flavor-singlet operators we adopt the shorthand notation $\bar{q}\Gamma q\equiv\bar{q}\Gamma t^{0}q$, with the factor $t^{0}$ always implied but suppressed. We use the notation $(x_{1},x_{2},x_{3},x_{4})\equiv(x,y,z,t)\equiv(\vec{x},t)$ interchangeably. With this notation, the $t$-correlator is defined as
$C_{\Gamma}(t)=\int d^{3}x\;\bigl\langle O_{\Gamma}^{a}(\vec{x},t)\,O_{\Gamma}^{a}(\mathbf{0},0)\bigr\rangle,$ (1)
and its spatial counterpart, the $z$-correlator, as
$C_{\Gamma}(z)=\int dx\,dy\,dt\;\bigl\langle O_{\Gamma}^{a}(x,y,z,t)\,O_{\Gamma}^{a}(\mathbf{0},0)\bigr\rangle.$ (2)
If a symmetry is exact, the correlators of two partners $A$ and $B$ satisfy $C_{A}(t)=C_{B}(t)$ for any $t$ and $C_{A}(z)=C_{B}(z)$ for any $z$.
Composite operators such as $O_{\Gamma}^{a}(x)$ require regularization and renormalization. Both $C_{\Gamma}(t)$ and $C_{\Gamma}(z)$ contain short-distance singularities; the leading divergence behaves as $t^{-3}$ or $z^{-3}$ in the continuum limit as the two operators approach coincidence, corresponding to a contact term $\sim a^{-3}$ on the lattice whose precise form is regularization-dependent. In what follows we concentrate on the $t$-correlator; the extension to the $z$-correlator is straightforward.
2.2 Renormalization and RG invariance
To quantify the deviation from degeneracy at a specific Euclidean time one may define a pointwise ratio
$\kappa_{AB}(t)=\frac{C_{A}(t)-C_{B}(t)}{C_{A}(t)+C_{B}(t)},\qquad t\neq 0.$ (3)
An analogous ratio $\kappa_{AB}(z)$ for spatial correlators has been employed to study $U(1)_{A}$ and $SU(2)_{L}\times SU(2)_{R}$ symmetry patterns in thermal QCD with optimal domain-wall fermions Chiu:2024jyz ; Chiu:2024bqx .
For lattice formulations that preserve chiral symmetry (e.g., overlap or domain-wall fermions) the renormalization constants of symmetry-related operators are identical, $Z_{A}=Z_{B}$, because the transformation that relates $O_{A}^{a}$ and $O_{B}^{a}$ is a symmetry of the regularized lattice action. Concrete examples in two-flavor QCD include:
  • under $U(1)_{A}$: the nonsinglet scalar ($S$, e.g. $\delta^{+}\sim\bar{d}u$) and pseudoscalar ($P$, e.g. $\pi^{+}\sim\bar{d}\gamma_{5}u$) are partners, hence $Z_{S}=Z_{P}$;
  • under $SU(2)_{L}\times SU(2)_{R}$: the nonsinglet vector ($V$, e.g. $\rho^{+}\sim\bar{d}\gamma_{i}u$) and axial-vector ($A$, e.g. $a_{1}^{+}\sim\bar{d}\gamma_{5}\gamma_{i}u$) are partners, hence $Z_{V}=Z_{A}$;
  • under $U(1)_{A}$ (tensor channel): the tensor ($T$, e.g. $\rho_{T}^{+}\sim\bar{d}\gamma_{4}\gamma_{i}u$) and axial-tensor ($X$, e.g. $b_{1}^{+}\sim\bar{d}\gamma_{5}\gamma_{4}\gamma_{i}u$) are partners, hence $Z_{T}=Z_{X}$.
Consequently, the ratio $\kappa_{AB}(t)$ in eq. (3) is renormalization-group invariant, as the common factor $Z_{A}^{2}=Z_{B}^{2}$ cancels between numerator and denominator.
2.3 Integrated spectral weights and the symmetry strength parameter
Although $\kappa_{AB}(t)$ is a valid probe, its dependence on $t$ and the statistical fluctuations at individual times make a channel-wide comparison of symmetry breaking cumbersome. A more robust global measure is obtained by integrating over Euclidean time, which sums the spectral weight in each channel. We therefore define the regularized susceptibility for channel $A$. The contact term at $t=0$ is subtracted to remove the short-distance singularity; for flavor-singlet and Cartan nonsinglet channels with scalar or tensor Dirac structure (for which the vacuum expectation value $\langle\bar{q}\Gamma_{A}t^{a}q\rangle$ is nonzero), the vacuum contribution is also subtracted. The integration runs over $t\in(0,1/T)$, where $T$ is the temperature, with the $t=0$ contact term excluded:
$\chi_{A}=\int_{0<t<1/T}dt\left[C_{A}(t)-\int d^{3}x\langle\bar{q}\Gamma_{A}t^{a}q\rangle_{\text{vev}}^{2}\right],$ (4)
where
$\langle\bar{q}\Gamma_{A}t^{a}q\rangle_{\text{vev}}=\frac{1}{V}\int d^{4}x\langle\bar{q}(x)\Gamma_{A}t^{a}q(x)\rangle_{\text{F}},\qquad V=\int d^{4}x.$ (5)
The subtraction of $C_{A}(0)$ and the vacuum term removes the dominant power divergence, leaving a multiplicatively renormalizable quantity: $\chi_{A}^{R}=Z_{A}^{2}\chi_{A}$. An analogous definition holds for channel $B$.
Because $Z_{A}=Z_{B}$ for symmetry partners, we can now construct a global, RG-invariant measure of symmetry breaking, the symmetry strength parameter:
$\boxed{\;\kappa_{AB}=\frac{\chi_{A}-\chi_{B}}{\chi_{A}+\chi_{B}},\;}$ (6)
where $\chi_{A}\geq\chi_{B}>0$ (with the convention $\chi_{A}\geq\chi_{B}$ adopted without loss of generality), and $0\leq\kappa_{AB}\leq 1$. This parameter provides an intuitive scale: $\kappa_{AB}=0$ indicates exact degeneracy of the integrated spectral weights (the symmetry is manifest), whereas $\kappa_{AB}\to 1$ signals maximal asymmetry. The construction is general and can be applied to study any symmetry in quantum field theory for which partner operators can be defined.
Remark on terminology. Throughout this work, “restoration” of a symmetry always refers to effective restoration — i.e., the corresponding RG‑invariant ratio $\kappa_{AB}$ becomes consistent with zero within statistical uncertainties. Exact restoration, which would require vanishing quark masses for $SU(2)_{L}\times SU(2)_{R}$ and additionally the complete suppression of the $U(1)_{A}$ anomaly for $U(1)_{A}$ restoration, is neither achievable in physical QCD nor required for the validity of our conclusions. This operational definition is implicit in all subsequent uses of “degeneracy”, “restoration”, and $\kappa_{AB}=0$”.
In the remainder of this paper we employ $\kappa_{AB}$ to investigate the restoration patterns of the $U(1)_{A}$ and $SU(2)_{L}\times SU(2)_{R}$ symmetries in high-temperature lattice QCD.
2.4 Symmetry Channels Involving Flavor Singlet Operators
The full chiral $SU(2)_{L}\times SU(2)_{R}$ multiplet for $J=0$ mesons includes not only the connected channels studied in this work, but also symmetry pairs involving flavor-singlet operators. In principle, each pair could be probed by the symmetry ratio $\kappa_{AB}$ defined in eq. (6), offering additional insight into chiral restoration.
For $N_{f}=2$ massless QCD, effective restoration of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry implies equalities among regularized susceptibilities as shown in eqs. (7)–(11),
$\chi^{s}_{S}$ $=\chi^{ns}_{P}\Longleftrightarrow\chi_{\sigma}=\chi_{\pi},$ (7)
$\chi^{s}_{P}$ $=\chi^{ns}_{S}\Longleftrightarrow\chi_{\eta}=\chi_{\delta},$ (8)
$\chi^{s}_{T_{k}}$ $=\chi^{ns}_{X_{k}}\Longleftrightarrow\chi_{\omega_{T}}=\chi_{b_{1}},\qquad k=1,2,3,$ (9)
$\chi^{s}_{X_{k}}$ $=\chi^{ns}_{T_{k}}\Longleftrightarrow\chi_{h_{1T}}=\chi_{\rho_{T}},\qquad k=1,2,3,$ (10)
$\chi^{ns}_{V_{k}}$ $=\chi^{ns}_{A_{k}}\Longleftrightarrow\chi_{\rho}=\chi_{a_{1}},\qquad k=1,2,3,$ (11)
where the superscripts $s$ and $ns$ denote singlet and nonsinglet channels, respectively. Recall that the physical $\eta^{\prime}$ corresponds to the $\eta$ singlet in $N_{f}=2$ QCD. The notations of meson operators and their corresponding regularized susceptibilies are summarized in appendix A.
On the other hand, effective restoration of $U(1)_{A}$ axial symmetry implies eqs. (12)–(15).
$\chi^{ns}_{S}$ $=\chi^{ns}_{P}\Longleftrightarrow\chi_{\delta}=\chi_{\pi},$ (12)
$\chi^{s}_{S}$ $=\chi^{s}_{P}\Longleftrightarrow\chi_{\sigma}=\chi_{\eta},$ (13)
$\chi^{ns}_{T_{k}}$ $=\chi^{ns}_{X_{k}}\Longleftrightarrow\chi_{\rho_{T}}=\chi_{b_{1}},\qquad k=1,2,3.$ (14)
$\chi^{s}_{T_{k}}$ $=\chi^{s}_{X_{k}}\Longleftrightarrow\chi_{\omega_{T}}=\chi_{h_{1T}},\qquad k=1,2,3,$ (15)
Each equality corresponds to a distinct RG-invariant symmetry ratio $\kappa_{AB}$ as defined in eq. (6).
Equality of renormalization constants from symmetry. For chirally symmetric fermion formulations, the symmetry relations in eqs. (7)–(15) imply specific equalities among renormalization constants.
For scalar and pseudoscalar channels, the partner relations yield:
  • From $SU(2)_{L}\times SU(2)_{R}$: $(\sigma,\pi)$ and $(\eta,\delta)$ are partners, giving $Z_{S}^{s}=Z_{P}^{ns}$ and $Z_{P}^{s}=Z_{S}^{ns}$.
  • From $U(1)_{A}$: $(\delta,\pi)$ and $(\sigma,\eta)$ are partners, giving $Z_{S}^{ns}=Z_{P}^{ns}$ and $Z_{S}^{s}=Z_{P}^{s}$.
Combining these gives:
$Z_{S}^{ns}=Z_{P}^{ns}=Z_{S}^{s}=Z_{P}^{s}.$
For tensor and axial-tensor channels, the same logic applies:
  • Under $SU(2)_{L}\times SU(2)_{R}$: $(T^{s},X^{ns})$ and $(X^{s},T^{ns})$ are partners, giving $Z_{T}^{s}=Z_{X}^{ns}$ and $Z_{X}^{s}=Z_{T}^{ns}$.
  • Under $U(1)_{A}$: $(T^{s},X^{s})$ and $(T^{ns},X^{ns})$ are partners, giving $Z_{T}^{s}=Z_{X}^{s}$ and $Z_{T}^{ns}=Z_{X}^{ns}$.
Combining these gives:
$Z_{T}^{ns}=Z_{X}^{ns}=Z_{T}^{s}=Z_{X}^{s}.$
For vector and axial-vector channels, the pattern is different:
  • Under $SU(2)_{L}\times SU(2)_{R}$, the nonsinglet vector $\rho$ and axial-vector $a_{1}$ are partners, giving $Z_{V}^{ns}=Z_{A}^{ns}$. Both currents are conserved for all $m_{q}$, so $Z_{V}^{ns}=Z_{A}^{ns}=1$.
  • The singlet vector $\omega\sim\bar{q}\gamma_{\mu}q$ is associated with the exact $U(1)_{V}$ symmetry of the QCD action, which is preserved for all $m_{q}$, so $Z_{V}^{s}=1$.
  • The singlet axial-vector $h_{1\mu}\sim\bar{q}\gamma_{5}\gamma_{\mu}q$ is not conserved due to the $U(1)_{A}$ anomaly, regardless of quark mass. Its divergence is $\partial^{\mu}h_{1\mu}=2m_{q}\bar{q}\gamma_{5}q+\frac{g^{2}}{16\pi^{2}}F\tilde{F}$. Unlike the conserved currents above, $h_{1\mu}$ is not protected by a conserved charge, so its renormalization constant $Z_{A}^{s}$ is not fixed to unity and must be determined non-perturbatively. The current itself renormalizes multiplicatively (without operator mixing), but $Z_{A}^{s}\neq Z_{V}^{s}$ and $Z_{A}^{s}\neq 1$ in general.
These relations ensure that, after appropriate contact term and vacuum subtractions, any ratio $\kappa_{AB}$ involving operators from the same renormalization constant multiplet is RG-invariant. This includes the nonsinglet ratios $\kappa_{PS}$, $\kappa_{VA}$, $\kappa_{TX}$ studied in this work, as well as singlet-involved ratios such as $\kappa_{\sigma,\pi}$, $\kappa_{\eta,\delta}$, $\kappa_{\omega_{T},b_{1}}$, $\kappa_{h_{1T},\rho_{T}}$, and $\kappa_{\omega_{T},h_{1T}}$. Consequently, ratios like $\kappa_{\sigma,\pi}$ and $\kappa_{\eta,\delta}$ are formally RG-invariant, just like their nonsinglet counterparts.
Despite this formal RG invariance, several practical obstacles prevent a clean extraction of $\kappa_{AB}$ for channels involving flavor-singlet operators in the present study:
  1. 1.
    Disconnected diagrams and stochastic noise.
    Correlators of the singlet operators $\sigma$ and $\eta^{\prime}$ receive contributions from disconnected Wick contractions, which are computationally demanding and statistically noisy in lattice simulations and require extensive stochastic sampling. By contrast, their nonsinglet partners involve only connected diagrams. This asymmetry in diagrammatic structure leads to very different signal-to-noise ratios and systematic uncertainties in the estimation of the susceptibilities $\chi_{A}$ and $\chi_{B}$, compromising a direct comparison.
  2. 2.
    Regularization of scalar and pseudoscalar singlet susceptibilities.
    Following eq. (4), the regularized susceptibilities for singlet channels are defined by subtracting the power-divergent contact term. For the scalar singlet $\sigma\sim\bar{q}q$, this requires subtracting the vacuum expectation value $\langle\bar{q}q\rangle^{2}$ to remove the power divergence $\propto 1/a^{2}$. For the pseudoscalar singlet $\eta^{\prime}\sim\bar{q}\gamma_{5}q$, parity ensures $\langle\bar{q}\gamma_{5}q\rangle=0$ in the continuum, so no vacuum subtraction is needed. However, the unsubtracted bare susceptibility contains power-divergent mixing with the topological charge density $F\tilde{F}$ through coefficients $\sim 1/a^{n}$. The contact term subtraction in eq. (4) removes these power-divergent mixings as well. After this subtraction, the resulting regularized susceptibility $\chi_{\eta^{\prime}}^{\mathrm{sub}}$ is free of all power divergences and multiplicatively renormalizable with $Z_{P}^{s}=Z_{S}^{ns}$. Crucially, all power-divergent mixings, including any power-divergent mixing with the topological susceptibility $\chi_{t}$, are completely removed by these subtractions.
  3. 3.
    Finite mixing with topological susceptibility.
    Even after all power divergences are removed, the $U(1)_{A}$ anomaly induces finite physical mixing between the singlet susceptibilities and the topological susceptibility $\chi_{t}$. For the scalar singlet, this mixing occurs through a scale-independent coefficient:
    $\chi_{\sigma}^{R}=Z_{S}^{2}\chi_{\sigma}^{\mathrm{sub}}+c_{S}\,\chi_{t}^{R},$
    where $c_{S}$ is a finite, scheme-dependent constant; its scale independence follows from the fact that $\chi_{\sigma}$ and $\chi_{t}$ share the same $CP$ quantum numbers and their mixing is mediated by a dimension-four operator whose anomalous dimension vanishes at leading order. For the pseudoscalar singlet, the mixing is through a scale-dependent coefficient:
    $\chi_{\eta^{\prime}}^{R}=Z_{S}^{2}\chi_{\eta^{\prime}}^{\mathrm{sub}}+\tilde{c}(\mu)\,\chi_{t}^{R},$
    with $\tilde{c}(\mu)$ evolving under the renormalization group due to the interplay with the axial anomaly Ward identity. These mixing coefficients are finite — they are not power-divergent and cannot be removed by the subtractions of eq. (4). Obtaining the fully renormalized susceptibilities $\chi_{\sigma}^{R}$ and $\chi_{\eta^{\prime}}^{R}$ therefore requires a non-perturbative determination of $c_{S}$ and $\tilde{c}(\mu)$ Martinelli:1995ty .
    Constructing fully RG-invariant ratios such as $\kappa_{\sigma,\pi}$ and $\kappa_{\eta,\delta}$ would then require performing a non-perturbative subtraction of the finite mixing terms. We defer this substantial task to future work.
    For overlap fermions with exact chiral symmetry, the topological charge can be defined via the index theorem ($Q_{t}=n_{+}-n_{-}$) through zero mode counting. In this case, $\chi_{t}$ requires no renormalization and the bare susceptibility equals the renormalized one: $\chi_{t}^{\mathrm{bare}}=\chi_{t}^{R}$. This simplifies the analysis, but the challenge of determining the finite mixing coefficients remains. In practice, at finite lattice spacing, gauge field dislocations can produce spurious near-zero modes that contaminate the zero-mode count; sufficiently smooth configurations are required for the index theorem identification to be unambiguous.
  4. 4.
    Tensor singlet channels: a cleaner probe.
    In contrast to the scalar and pseudoscalar singlets, the tensor singlet operators $\omega_{T,k}\equiv T_{k}^{s}=\bar{q}\gamma_{4}\gamma_{k}q$ and $h_{1T,k}\equiv X_{k}^{s}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ avoid both power-divergent mixings and finite mixing with $\chi_{t}$. Their renormalization is purely multiplicative with $Z_{T}^{s}=Z_{X}^{s}$ in chirally symmetric formulations, making the ratio $\kappa_{\omega_{T},h_{1T}}=(\chi_{\omega_{T}}-\chi_{h_{1T}})/(\chi_{\omega_{T}}+\chi_{h_{1T}})$ RG-invariant and theoretically clean. Mixed ratios $\kappa(\chi_{\omega_{T}},\chi_{b_{1}})$ and $\kappa(\chi_{h_{1T}},\chi_{\rho_{T}})$ similarly probe $SU(2)_{L}\times SU(2)_{R}$ restoration between singlet and nonsinglet tensor partners. Future studies with improved techniques for disconnected diagrams may enable reliable investigation of these channels.
Given these considerations, the present study focuses initially on the connected, nonsinglet channels where the RG‑invariant ratio $\kappa_{AB}$ rests on the firmest theoretical and numerical ground:
  • $\kappa_{PS}$ ($\delta$-$\pi$) for $U(1)_{A}$,
  • $\kappa_{VA}$ ($\rho$-$a_{1}$) for $SU(2)_{L}\times SU(2)_{R}$,
  • $\kappa_{TX}$ ($\rho_{T}$-$b_{1}$) for $U(1)_{A}$ through a different Dirac structure.
These choices provide a clean, theoretically unambiguous, and statistically precise set of observables. The consistent picture emerging from these three channels (presented in section 4 and section 5) gives strong evidence for the relative restoration scales of $U(1)_{A}$ and $SU(2)_{L}\times SU(2)_{R}$ symmetries.
3 Lattice setup
We generate gauge ensembles using hybrid Monte Carlo (HMC) simulations of lattice QCD with $N_{f}=2+1+1$ optimal domain-wall quarks Chiu:2002ir ; Chiu:2015sea at the physical point. The simulations are performed on $32^{3}\times(16,12,10,8)$ lattices with the plaquette gauge action Wilson:1974sk at three values of $\beta=6/g^{2}=(6.15,6.18,6.20)$, corresponding to lattice spacings $a\simeq(0.075,0.069,0.064)$ fm. These ensembles are produced with the same actions Chiu:2011bm ; Chen:2014hyy and algorithms as their counterparts on larger $64^{3}\times(20,16,12,10,8,6)$ lattices Chen:2022fid , but at one-eighth of the spatial volume. Simulations are carried out on GPU clusters equipped with various NVIDIA GPUs.
After initial thermalization, gauge configurations are sampled and distributed among 16–32 independent simulation units, each performing a separate HMC stream. In each stream, one configuration is sampled every five trajectories. All sampled configurations from all streams are combined to obtain the final ensemble. Lattice parameters and statistics for the meson $t$-correlator calculations are listed in table 1. The temperatures covered range from $\sim 164$ to $385$ MeV, all above the pseudocritical temperature $T_{c}\sim 150$ MeV.
$\beta$ aa[fm] $N_{x}$ $N_{t}$ TT[MeV] $N_{\rm confs}$ $(m_{u/d}a)_{\rm res}$ $(m_{s}a)_{\rm res}$ $(m_{c}a)_{\rm res}$
6.15 0.075 32 16 164 359 $6.2(5){\times}10^{-5}$ $3.1(4){\times}10^{-5}$ $1.1(7){\times}10^{-5}$
6.18 0.069 32 16 179 324 $4.9(7){\times}10^{-5}$ $5.8(1.7){\times}10^{-5}$ $2.7(9){\times}10^{-5}$
6.20 0.064 32 16 192 588 $3.5(4){\times}10^{-5}$ $2.3(2){\times}10^{-5}$ $6.5(1.1){\times}10^{-6}$
6.15 0.075 32 12 219 409 $9.0(1.4){\times}10^{-5}$ $6.0(8){\times}10^{-5}$ $1.5(3){\times}10^{-6}$
6.18 0.069 32 12 238 781 $1.9(2){\times}10^{-5}$ $1.6(1){\times}10^{-5}$ $3.8(5){\times}10^{-6}$
6.20 0.064 32 12 257 514 $1.8(6){\times}10^{-5}$ $1.6(6){\times}10^{-5}$ $9.4(4.2){\times}10^{-6}$
6.15 0.075 32 10 263 496 $2.4(4){\times}10^{-5}$ $2.0(3){\times}10^{-5}$ $7.5(2.6){\times}10^{-6}$
6.18 0.069 32 10 286 377 $4(8){\times}10^{-5} 2.2$$2(7){\times}10^{-5}$ $9.7(3.9){\times}10^{-6}$
6.20 0.064 32 10 308 481 $5.8(1.5){\times}10^{-6}$ $5.1(1.2){\times}10^{-6}$ $1.4(2){\times}10^{-6}$
6.15 0.075 32 8 328 640 $3.5(8){\times}10^{-5}$ $2.9(6){\times}10^{-5}$ $1.2(3){\times}10^{-5}$
6.18 0.069 32 8 357 302 $1.3(7){\times}10^{-5}$ $1.2(6){\times}10^{-5}$ $4.9(2.2){\times}10^{-6}$
6.20 0.064 32 8 385 468 $6.3(1.9){\times}10^{-6}$ $6.0(1.8){\times}10^{-6}$ $3.0(9){\times}10^{-6}$
Table 1: Lattice parameters and statistics of the twelve gauge ensembles used in this work. The last three columns give the residual masses of $u/d$, $s$, and $c$ quarks Chen:2012jya .
Lattice spacings and quark masses ($u/d$, $s$, $c$) are determined on $32^{3}\times 64$ lattices with $\{460,636,726\}$ configurations for $\beta=\{6.15,6.18,6.20\}$, respectively. The lattice spacing is fixed using the Wilson flow Narayanan:2006rf ; Luscher:2010iy with the condition $\{t^{2}\langle E(t)\rangle\}|_{t=t_{0}}=0.3$ and input $\sqrt{t_{0}}=0.1416(8)$ fm Bazavov:2015yea . The resulting spacings are listed in table 2. Physical quark masses are obtained by tuning the bare masses so that the lowest-lying states extracted from time-correlation functions of the meson operators $\{\bar{u}\gamma_{5}d,\bar{s}\gamma_{i}s,\bar{c}\gamma_{i}c\}$ agree with the physical masses of $\pi^{\pm}(140)$, $\phi(1020)$, and $J/\psi(3097)$. The tuned bare quark masses are also given in table 2.
$\beta$ aa[fm] $m_{u/d}a$ $m_{s}a$ $m_{c}a$
6.15 0.0751(5) 0.00200 0.064 0.705
6.18 0.0690(5) 0.00180 0.058 0.626
6.20 0.0641(4) 0.00125 0.040 0.550
Table 2: Lattice spacings and bare quark masses for $N_{f}=2+1+1$ lattice QCD with optimal domain-wall quarks at the physical point.
[…]
We now summarize the notations and conventions used in this work.
The correlation function of off-diagonal flavor-nonsinglet meson interpolator $\bar{q}_{1}\Gamma q_{2}$ (e.g., $\bar{u}\Gamma d$) on a lattice with $(N_{x},N_{y},N_{z},N_{t})$ sites is computed as
$C_{\Gamma}(x)=\bigl\langle(\bar{q}_{1}\Gamma q_{2})_{x}(\bar{q}_{1}\Gamma q_{2})_{0}^{\dagger}\bigr\rangle=\Bigl\langle\operatorname{tr}\bigl[\Gamma(D_{c}+m_{1})^{-1}_{0,x}\,\Gamma(D_{c}+m_{2})^{-1}_{x,0}\bigr]\Bigr\rangle_{\text{confs}},$ (16)
where $(D_{c}+m_{q})^{-1}$ denotes the valence quark propagator with mass $m_{q}$ in lattice QCD with exact chiral symmetry Chiu:1998eu , $\operatorname{tr}$ is the trace over color and Dirac indices, and $\langle\cdots\rangle_{\text{confs}}$ denotes the average over gauge configurations. Here $x=(x_{1},x_{2},x_{3},x_{4})=(x,y,z,t)$; an overall sign arising from $\gamma_{4}\Gamma^{\dagger}\gamma_{4}=\pm\Gamma$ has been suppressed. The temporal correlator is defined as
$C_{\Gamma}(t,T)=\sum_{x_{1},x_{2},x_{3}}C_{\Gamma}(x),$ (17)
where $T=1/(N_{t}a)$ is the temperature. The regularized susceptibility (4) can be written as
$\chi_{\Gamma}(T)=\sum_{t=1}^{N_{t}-1}C_{\Gamma}(t,T).$ (18)
In this study we focus on the off-diagonal flavor-nonsinglet operator $\bar{u}\Gamma d/\sqrt{2}$, with
$\Gamma=\{\mbox{1\hskip-3.41432ptI},\gamma_{5},\gamma_{k},\gamma_{5}\gamma_{k},\gamma_{4}\gamma_{k},\gamma_{5}\gamma_{4}\gamma_{k},k=1,2,3\},$
corresponding to the scalar (S), pseudoscalar (P), vector (V), axial-vector (A), tensor-vector (T), and axial-tensor-vector (X) channels, respectively.
Thanks to $S_{3}$ symmetry, the correlators satisfy $C_{V_{1}}=C_{V_{2}}=C_{V_{3}}$, $C_{A_{1}}=C_{A_{2}}=C_{A_{3}}$, $C_{T_{1}}=C_{T_{2}}=C_{T_{3}}$, and $C_{X_{1}}=C_{X_{2}}=C_{X_{3}}$. To improve statistics, we average over the three spatial components for each channel, e.g.,
$C_{V}(t,T)=\frac{1}{3}\sum_{k=1}^{3}C_{V_{k}}(t,T),$
and similarly for $A$, $T$ and $X$.
These averaged correlators $C_{\Gamma}(t,T)$ are used to compute the regularized susceptibility of off-diagonal flavor nonsinglet mesons,
$\chi_{\Gamma}(T)=\sum_{t=1}^{N_{t}-1}C_{\Gamma}(t,T)$
and the RG-invariant symmetry ratios $\kappa_{AB}$ defined in (6) for the channel pairs $VA$, PSPS, and $TX$. Here $\kappa_{VA}$ probes $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry restoration, while $\kappa_{PS}$ and $\kappa_{TX}$ probe $U(1)_{A}$ axial symmetry restoration.
4 Results
We begin by analyzing the temporal correlators $C_{\Gamma}(t)$ of the $\bar{u}\Gamma d$ bilinears at the three lowest temperatures, $T=(164,179,193)$ MeV, displayed in figure 1. Our focus is on the degeneracy patterns among the $(V,A)$, $(P,S)$, and $(T,X)$ channels, which respectively reflect the restoration of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry and $U(1)_{A}$ axial symmetry.
As shown in figure 1, the correlators follow the ordering
$C_{P}(t)\gtrsim C_{S}(t)>C_{V}(t)\gtrsim C_{A}(t)>C_{T}(t)\gtrsim C_{X}(t)$
at all temperatures. This hierarchy corresponds to the ordering of meson thermal masses,
$m_{P}\lesssim m_{S}<m_{V}\lesssim m_{A}<m_{T}\lesssim m_{X},$
which remains consistent across the full temperature range studied (164–385 MeV).
Figure 1: $t$-correlators of $\bar{u}\Gamma d$ at the three lowest temperatures. The dashed lines connecting the data points in each channel are shown only to guide the eye.
Figure 2: Regularized susceptibilities of $\bar{u}\Gamma d$ for three lattice spacings aa=(0.075, 0.069, 0.064) fm and twelve temperatures from 164-385 MeV. The dashed lines connecting the data points in each channel are shown only to guide the eye.
Within the resolution of our data, no splitting is observed inside the $(P,S)$, $(V,A)$, or $(T,X)$ doublets, except for a mild deviation in the $(P,S)$ channel at $T=164$ MeV on the coarsest lattice ($a=0.075$ fm), shown in the upper panel of figure 1. Since degeneracy in the $(P,S)$ and $(T,X)$ channels signals $U(1)_{A}$ restoration, the observed discrepancy is likely a lattice artifact due to finite lattice spacing.
To cleanly address this issue in the continuum limit, a renormalization-group (RG) invariant measure of degeneracy is required—one that can be extrapolated to $a\to 0$ and allows quantitative comparison of symmetry breaking across different channels. Standard probes, such as thermal hadron masses or single-time correlator values, can be ambiguous or sensitive to analysis details. Instead, we employ the RG-invariant symmetry ratio $\kappa_{AB}$ defined in Eq. (6), which integrates spectral information, provides clear normalization, and enables direct comparison between $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ breaking in all channels.
Accordingly, we compute the regularized susceptibilities $\chi_{P}$, $\chi_{S}$, $\chi_{V}$, $\chi_{A}$, $\chi_{T}$, and $\chi_{X}$, shown in figure 2, as well as the RG-invariant symmetry ratios $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ presented in figure 3. These quantities are obtained for three lattice spacings $a=(0.075,0.069,0.064)$ fm and twelve temperatures from 164 to 385 MeV.
Numerical values of the correlators, susceptibilities, and symmetry ratios are provided in appendix B.
In figure 2, the regularized susceptibilities show no splitting within any doublet, again except for the $(P,S)$ channel at $T=164$ MeV on the coarsest lattice. This aligns with the behavior seen in the correlators and reinforces the interpretation that the observed $(P,S)$ deviations at the lowest temperature are artifacts of finite lattice spacing.
From figure 3, a clear hierarchy emerges:
$\kappa_{PS}\gg\kappa_{TX}\simeq\kappa_{VA},$
which holds for all three lattice spacings and across the entire temperature range. Note that the vertical scale for $\kappa_{PS}$ (upper panel) is two orders of magnitude larger than those for $\kappa_{VA}$ and $\kappa_{TX}$ (lower panels). If one were to use $\kappa_{PS}$ and $\kappa_{VA}$ at finite $a$ as measures of $U(1)_{A}$ and $SU(2)_{L}\times SU(2)_{R}$ breaking, respectively, the data would suggest that $U(1)_{A}$ restoration occurs at a higher temperature than chiral symmetry restoration. Conversely, using $\kappa_{TX}$ as the $U(1)_{A}$ indicator would imply that both symmetries are restored at nearly the same temperature, even at finite spacing. This apparent contradiction can only be resolved by taking the continuum limit, which we address in the next section.
5 Continuum limit
After testing the data for $\kappa_{PS}(a,T)$, $\kappa_{VA}(a,T)$, and $\kappa_{TX}(a,T)$ against various models, we find that at fixed lattice spacing $a$, the temperature dependence of each $\kappa$ can be well described by a simple power law:
$\kappa(a,T)=A(a)\bigl(T[\text{GeV}]\bigr)^{-B(a)},$ (19)
where the temperature $T$ is expressed in GeV.
For each lattice spacing, we first extract the parameters $A(a)$ and $B(a)$ via a log–linear fit:
$\ln\kappa(a,T)=\ln A(a)-B(a)\ln\left(T[\text{GeV}]\right).$
The resulting fitted parameters $A(a)$ and $B(a)$, together with the corresponding $\chi^{2}$/dof, are summarized for $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ in the first three rows of table 3.
Figure 3: RG-invariant symmetry ratios $\kappa_{VA}$ and $\{\kappa_{PS},\kappa_{TX}\}$ for the $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ chiral symmetries of $(u,d)$ quarks, computed at three lattice spacings $a$ = (0.075, 0.069, 0.064) fm and twelve temperatures in the range 164-385 MeV. The solid line and its band denote the continuum-extrapolated value and its uncertainty from 2D global fit.
In the second step, we extrapolate $A(a)$ and $B(a)$ to the continuum limit ($a\to 0$). We consider both linear and exponential extrapolations in $a^{2}$:
$f(a)=f_{0}+f_{1}a^{2},\quad\text{or}\quad f(a)=f_{0}\exp(f_{1}a^{2}),$
and choose between them based on the behavior of the data. If a linear fit yields an unphysical $f_{0}<0$, we adopt the exponential form, which includes higher-order contributions in $a^{2}$ and ensures positivity of $f_{0}$. For cases where both forms are viable, we select the one that gives the better fit to the data. The requirement $f_{0}>0$ is physically motivated: $A_{0}<0$ would imply $\kappa(a=0,T)<0$, contradicting the definition $\kappa_{AB}\geq 0$ from Eq. (6), while $B_{0}<0$ would imply that $\kappa$ increases with $T$, inconsistent with the expected decrease of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ breakings as temperature rises.
The continuum-extrapolated parameters $A_{0}$ and $B_{0}$, along with their $\chi^{2}$/dof and the chosen fitting functions, are listed for $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ in the last two rows of table 3.
$a$ [fm] $\kappa_{PS}$ $\kappa_{VA}$ $\kappa_{TX}$
$A\times 10^{7}$ $B$ $\chi^{2}$/dof $A\times 10^{8}$ $B$ $\chi^{2}$/dof $A\times 10^{8}$ $B$ $\chi^{2}$/dof
0.075 $2.53(2.08)$ 5.33(56) 0.13 $3.38(1.03)$ 3.63(20) 0.54 $3.85(3.63)$ 3.65(69) 0.28
0.069 $2.19(1.92)$ 4.55(67) 0.13 $3.34(89)$ 3.41(20) 1.45 $7.43(3.62)$ 2.92(36) 1.31
0.064 $1.04(83)$ 4.59(61) 0.34 $3.29(37)$ 3.18(10) 0.26 $8.21(1.24)$ 2.36(14) 0.79
cont. extrap. exp. linear $\chi^{2}$/dof linear linear $\chi^{2}$/dof linear exp. $\chi^{2}$/dof
2-step 0.14(0.46) 2.64(2.42) 0.10, 0.25 3.08(2.60) 2.12(58) <0.1<0.1, <0.1<0.1 17.66(9.17) 0.82(0.36) 0.10, <0.1<0.1
2D 0.60(2.28) 1.47(2.79) 0.35 2.71(2.17) 2.20(48) 0.56 26.29(4.43) 0.47(0.11) 0.75
Table 3: Fitted parameters $A(a)$, $B(a)$, and $\chi^{2}$/dof for the power-law ansatz $\kappa(a,T)=A(a)(T[\text{GeV}])^{-B(a)}$ for $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ at three lattice spacings. The fourth row indicates the functional form (linear or exponential) used in the continuum extrapolation of $A(a)$ and $B(a)$ for the two-step fit and the two-dimensional global fit. The last two rows present the continuum results ($a=0$) obtained via the two-step and 2D global fitting methods.
Guided by the functional forms adopted in the two-step method (fourth row of table 3), we also perform a simultaneous two-dimensional (2D) global fit to all 12 data points of each $\kappa_{AB}(a,T)$ using one of the following models:
$\kappa(a,T)$ $=(A_{0}+A_{1}a^{2})\bigl(T[\text{GeV}]\bigr)^{-(B_{0}+B_{1}a^{2})},$ (20)
$\kappa(a,T)$ $=A_{0}\exp(A_{1}a^{2})\bigl(T[\text{GeV}]\bigr)^{-(B_{0}+B_{1}a^{2})},$ (21)
$\kappa(a,T)$ $=(A_{0}+A_{1}a^{2})\bigl(T[\text{GeV}]\bigr)^{-B_{0}\exp(B_{1}a^{2})},$ (22)
where Eq. (21) is used for $\kappa_{PS}(a,T)$, Eq. (20) for $\kappa_{VA}(a,T)$, and Eq. (22) for $\kappa_{TX}(a,T)$.
The resulting continuum parameters $A_{0}$ and $B_{0}$ and the corresponding $\chi^{2}$/dof are given in the last row of table 3. The values obtained from the 2D global fit are consistent with those from the two-step method.
Statistically, the 2D-fit is more reliable because it performs a simultaneous minimization over the entire dataset, which ensures a globally optimized balance between parameters $A$ and $B$, and generally yields a better $\chi^{2}$ per degree of freedom. In contrast, the 2-step fit first fits each $a$-group independently to obtain local estimates $A(a)$ and $B(a)$, then performs a secondary regression on those results. This "fit-of-fits" approach propagates errors less efficiently and can introduce bias if a particular $a$-group exhibits larger statistical fluctuations. Such behavior is evident in table 3, where the $\chi^{2}$/d.o.f. values of $\kappa_{VA}$ and $\kappa_{TX}$ at $a=0.069$ fm are significantly larger than those at other lattice spacings.
We therefore adopt the 2D-fit results for $A_{0}$ and $B_{0}$ to obtain the continuum-extrapolated $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$. These are shown as solid curves in Fig. 3, with the error bands indicating the uncertainty. Note that the vertical scale for $\kappa_{PS}$ (upper panel) is two orders of magnitude larger than those for $\kappa_{VA}$ and $\kappa_{TX}$ (lower panels). The continuum extrapolated values of $\kappa_{PS}$, $\kappa_{VA}$ and $\kappa_{TX}$ are compatible with one another across the temperature range 150–400 MeV, within our precision. This is illustrated explicitly in table 4, which lists the continuum values at several representative temperatures.
$T[\text{MeV}]$ $\kappa_{PS}(a=0)$ $\kappa_{VA}(a=0)$ $\kappa_{TX}(a=0)$
150 $(9.7\pm 63.2)\times 10^{-7}$ $(1.77\pm 2.15)\times 10^{-6}$ $(6.42\pm 1.68)\times 10^{-7}$
164 $(8.5\pm 53.8)\times 10^{-7}$ $(1.46\pm 1.73)\times 10^{-6}$ $(6.16\pm 1.57)\times 10^{-7}$
179 $(7.5\pm 45.8)\times 10^{-7}$ $(1.20\pm 1.39)\times 10^{-6}$ $(5.91\pm 1.47)\times 10^{-7}$
219 $(5.6\pm 31.7)\times 10^{-7}$ $(7.69\pm 8.36)\times 10^{-7}$ $(5.37\pm 1.25)\times 10^{-7}$
257 $(4.4\pm 23.6)\times 10^{-7}$ $(5.41\pm 5.60)\times 10^{-7}$ $(4.98\pm 1.10)\times 10^{-7}$
385 $(2.4\pm 11.3)\times 10^{-7}$ $(2.22\pm 2.05)\times 10^{-7}$ $(4.12\pm 0.81)\times 10^{-7}$
Table 4: Continuum extrapolated values of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ at selected temperatures between 150 and 400 MeV.
6 Discussion and Outlook
In this work, we have introduced a renormalization-group invariant observable, the symmetry strength parameter $\kappa_{AB}$, designed to provide a quantitative and scheme-independent measure of symmetry breaking in QCD. Using this ratio, we have performed a systematic lattice study of the relative strength of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry breaking and $U(1)_{A}$ axial symmetry breaking in finite-temperature QCD with $N_{f}=2+1+1$ flavors, employing optimal domain-wall fermions at the physical point. Our analysis spans three lattice spacings and twelve temperatures in the range 164164385385 MeV, allowing for controlled continuum extrapolations across the chiral crossover region.
We examined three independent symmetry-breaking channels in the nonsinglet sector with connected correlators, namely: the scalar–pseudoscalar channel sensitive to $U(1)_{A}$ breaking ($\kappa_{PS}$), the vector–axial-vector channel probing chiral symmetry breaking ($\kappa_{VA}$), and an additional $U(1)_{A}$-sensitive tensor–axial-tensor channel ($\kappa_{TX}$). At finite lattice spacing, these channels exhibit a clear hierarchy: $\kappa_{PS}>\kappa_{TX}\sim\kappa_{VA}$. However, this hierarchy collapses in the continuum limit, where all three symmetry strength parameters become statistically indistinguishable within our current precision.
This degeneracy constitutes a robust, model-independent result obtained from a chirally symmetric lattice formulation and demonstrates that discretization effects play a central role in apparent differences among symmetry-breaking channels. From a physical perspective, our findings indicate that chiral and axial symmetry-breaking effects for the nonsinglet sector in QCD become comparably suppressed over a narrow temperature interval near the chiral crossover. We find no evidence for a parametrically separated restoration scale for $U(1)_{A}$ symmetry relative to $SU(2)_{L}\times SU(2)_{R}$ symmetry in the continuum limit—a result that places stringent quantitative constraints on phenomenological descriptions of finite-temperature QCD that rely on delayed axial symmetry restoration.
Within a renormalization-group framework, the observed continuum-limit degeneracy of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ implies that $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetry-breaking effects for connected correlators become comparably suppressed near $T_{c}$. This suggests that the corresponding symmetry-breaking operators acquire similar infrared relevance for nonsinglet observables. However, as we discuss in section 6.1, the full effective restoration pattern—including singlet channels—reveals a more intricate structure that refines simple RG scenarios based solely on nonsinglet data.
6.1 Connections to topological susceptibility and singlet sector
An important subtlety involves the disconnected parts of the scalar singlet ($\sigma$) and pseudoscalar singlet ($\eta$) mesons in $N_{f}=2$ QCD. Recall that the physical $\eta^{\prime}$ corresponds to the $\eta$ singlet in $N_{f}=2$ QCD. Upon effective restoration of full $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry, i.e., satisfaction of Eqs. (11), (7) and (8), the latter two equations involving both nonsinglet and singlet channels yield
$\chi_{\sigma}-\chi_{\delta}=\chi_{\pi}-\chi_{\delta}=\chi_{\pi}-\chi_{\eta},$ (23)
which relates the disconnected parts of $\chi_{\sigma}$ and $\chi_{\eta}$:
$\chi_{\text{disc}}\equiv\chi_{\sigma}-\chi_{\delta}=\chi_{\pi}-\chi_{\eta}\equiv\chi_{5,\text{disc}},$ (24)
where
$\chi_{\text{disc}}$ $=\frac{1}{V}\left\{\left\langle[\mathrm{Tr}(D_{c}+m_{q})^{-1}]^{2}\right\rangle-\left\langle\mathrm{Tr}(D_{c}+m_{q})^{-1}\right\rangle^{2}\right\},$ (25)
$\chi_{5,\text{disc}}$ $=\frac{1}{V}\left\{\left\langle[\mathrm{Tr}\gamma_{5}(D_{c}+m_{q})^{-1}]^{2}\right\rangle-\left\langle\mathrm{Tr}\gamma_{5}(D_{c}+m_{q})^{-1}\right\rangle^{2}\right\}.$ (26)
In lattice QCD with exact chiral symmetry, the topological charge $Q_{t}$ satisfies
$Q_{t}=m_{q}\mathrm{Tr}\gamma_{5}(D_{c}+m_{q})^{-1},\quad\forall m_{q}>0.$ (27)
Assuming Eq. (27) can be measured precisely (e.g., with overlap fermions via index theorem), Eq. (24) implies
$\chi_{\text{disc}}=\chi_{5,\text{disc}}=\frac{\chi_{t}}{m_{q}^{2}},$ (28)
where $\chi_{t}$ is the topological susceptibility,
$\chi_{t}=\frac{1}{V}\left(\langle Q_{t}^{2}\rangle-\langle Q_{t}\rangle^{2}\right).$ (29)
Thus, upon effective restoration of full $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry, $\chi_{t}/m_{q}^{2}$ equals any of the three differences: $\chi_{\sigma}-\chi_{\delta}$, $\chi_{\pi}-\chi_{\delta}$, or $\chi_{\pi}-\chi_{\eta}$. However, Eq. (23) presents a puzzle: how could $(\chi_{\pi}-\chi_{\delta})$, which involves only connected correlators, equal $\chi_{\text{disc}}=\chi_{t}/m_{q}^{2}$, which arises solely from disconnected diagrams? This puzzle is resolved by considering the effective restoration of full $U(1)_{A}$ axial symmetry, which requires both Eq. (12) (nonsinglet) and Eq. (13) (singlet) to be satisfied. Subtracting Eq. (12) from Eq. (13) gives
$\chi_{\sigma}-\chi_{\delta}=\chi_{\eta}-\chi_{\pi},$ (30)
which contradicts Eq. (23) unless all susceptibility differences in both equations vanish—implying $\chi_{t}=0$. In other words, full effective restoration of both $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries requires vanishing topological susceptibility.
This leads to a natural two-stage hierarchical effective restoration scenario illustrated in Fig. 4:
  • Stage 1 ($T\sim T_{c}^{ns}\simeq T_{c}\sim 156$ MeV): Effective restoration of both $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries in the nonsinglet sector, where connected correlators become degenerate:
    • $U(1)_{A}$: $\chi_{\pi}\approx\chi_{\delta}$ (Eq. 12), $\chi_{\rho_{T}}\approx\chi_{b_{1}}$ (Eq. 14)
    • $SU(2)_{L}\times SU(2)_{R}$: $\chi_{\rho}\approx\chi_{a_{1}}$ (Eq. 11)
  • Stage 2 ($T\sim T_{c}^{s}\gg T_{c}^{ns}$): Full effective restoration including singlet and mixed singlet-nonsinglet channels, requiring $\chi_{t}\to 0$, $\kappa_{\omega_{T},h_{1T}}\to 0$ and satisfaction of all relations (eqs. (7)–(15)) including eqs. (23) and (30):
    • $SU(2)_{L}\times SU(2)_{R}$: $\chi_{\sigma}=\chi_{\pi}$, $\chi_{\eta}=\chi_{\delta}$, $\chi_{\omega_{T}}=\chi_{b_{1}}$, $\chi_{h_{1T}}=\chi_{\rho_{T}}$
    • $U(1)_{A}$: $\chi_{\sigma}=\chi_{\eta}$, $\chi_{\omega_{T}}=\chi_{h_{1T}}$
Figure 4: Schematic illustration of the two-stage hierarchical restoration scenario for light $(u,d)$ quarks. Stage 1 (nonsinglet restoration) occurs around $T_{c}^{ns}\sim 156$ MeV, while Stage 2 (full effective restoration including singlets) requires $\chi_{t}\to 0$ and $\kappa_{\omega_{T},h_{1T}}\to 0$ at a much higher temperature $T_{c}^{s}\gg T_{c}^{ns}$. The $\kappa_{AB}$ ratios for nonsinglet channels ($\kappa_{PS}$, $\kappa_{VA}$, $\kappa_{TX}$) probe Stage 1, while tensor singlet and mixed singlet-nonsinglet tensor ratios ($\kappa_{\omega_{T},h_{1T}}$, $\kappa(\chi_{\omega_{T}},\chi_{b_{1}})$, $\kappa(\chi_{h_{1T}},\chi_{\rho_{T}})$) and direct $\chi_{t}$ measurements probe Stage 2.
In this picture, $\chi_{\pi}-\chi_{\delta}$ (probing nonsinglet $U(1)_{A}$ restoration) is distinct from $\chi_{\sigma}-\chi_{\delta}$ and $\chi_{\pi}-\chi_{\eta}$ (probing mixed singlet-nonsinglet channel restoration). The former can approach zero while the latter remain finite, reflecting the persistence of topological fluctuations. The key insight is that both $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries exhibit two-stage restoration: they first restore in the nonsinglet sector around $T_{c}^{ns}\sim 156$ MeV, and only later restore fully (including singlet channels) at a much higher temperature $T_{c}^{s}\gg T_{c}^{ns}$ when topological fluctuations are largely suppressed ($\chi_{t}\to 0$).
6.2 Hierarchical restoration of chiral and axial symmetries
Our analysis of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ provides direct evidence for Stage 1 (nonsinglet restoration) of this hierarchical restoration. The near‑vanishing of $\kappa_{PS}$ and $\kappa_{TX}$ around $T_{c}$ demonstrates that $U(1)_{A}$‑anomaly effects on connected correlators become negligible in this regime, signaling effective $U(1)_{A}$ restoration in the nonsinglet sector. Similarly, $\kappa_{VA}\to 0$ indicates effective $SU(2)_{L}\times SU(2)_{R}$ restoration in the nonsinglet sector.
[…]
For $N_{f}=2+1+1$ QCD at the physical point, the chiral transition is a smooth crossover. Consequently, $\chi_{\sigma}-\chi_{\delta}$ or $\chi_{\eta}-\chi_{\pi}$ never vanishes identically but gradually diminishes as $T$ increases above $T_{c}\sim T_{c}^{ns}$. The relation $\chi_{\sigma}^{R}-\chi_{\delta}^{R}=\chi_{t}/m_{q}^{2}$ is obscured in practice by lattice artifacts, residual chiral symmetry breaking, anomalous renormalization uncertainties, and the inherent nonzero difference in a crossover. Lattice determinations of $\chi_{t}(T)$ suffer from significant discretization artifacts, and a consistent continuum picture has not yet emerged across different studies (see e.g., Petreczky:2016vrs ; Borsanyi:2016ksw ; Chen:2022fid ; Athenodorou:2022aay ; Kotov:2025ilm ). In general, for $N_{f}=2+1$ and $N_{f}=2+1+1$ QCD at physical masses, $\chi_{t}(T)$ remains sizeable up to $T\gg T_{c}$, indicating that topological fluctuations are not fully suppressed until well above the chiral crossover.
Notably, in lattice QCD with exact chiral symmetry, $\chi_{t}(T)$ for physical $N_{f}=2+1$ and $N_{f}=2+1+1$ QCD can be much larger than that of $N_{f}=2$ QCD in the chiral limit (see appendix A of ref. Mao:2009sy ). This raises an intriguing possibility: in the $N_{f}=2$ chiral limit, the two-stage hierarchy might nearly collapse, with $T_{c}^{\mathrm{s}}\gtrsim T_{c}^{\mathrm{ns}}\sim T_{c}$, implying that full effective restoration (including singlet channels) could occur much closer to the chiral transition temperature. Our simulations at physical masses cannot address this scenario directly, as the strange and charm quarks explicitly break the symmetry and enhance topological fluctuations.
Our $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ measurements, being ratios, cancel many systematics and cleanly show the trend toward symmetry restoration in the nonsinglet sector. The persistence of $\chi_{t}>0$ for $T\gtrsim T_{c}$ in lattice studies is therefore not in tension with $\kappa_{PS},\kappa_{VA},\kappa_{TX}\to 0$; rather, it reflects the hierarchy $T_{c}^{ns}\approx T_{c}<T_{c}^{s}$. The anomaly thus influences different observables in distinct ways and on different temperature scales. While $\chi_{t}$ probes the global topological charge distribution and remains finite until much higher temperatures (falling off with a power-law consistent with dilute instanton gas prediction Gross:1980br , although the overall normalization may differ significantly), nonsinglet meson correlators are sensitive primarily to the anomaly-induced splitting within chiral partners—a splitting that can diminish due to thermal screening even while topological fluctuations themselves persist.
Probing full effective restoration with tensor singlet channels
While $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ cleanly signal nonsinglet restoration near $T_{c}^{ns}$, the singlet sector remains numerically challenging. As discussed in section 2.4, the tensor singlet operators $\omega_{T,k}\equiv T_{k}^{s}=\bar{q}\gamma_{4}\gamma_{k}q$ and $h_{1T,k}\equiv X_{k}^{s}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ offer a theoretically cleaner alternative to scalar/pseudoscalar singlets. For chirally symmetric fermions, their renormalization constants satisfy $Z_{T}^{s}=Z_{X}^{s}$, and after contact term and vacuum subtraction, their susceptibilities are multiplicatively renormalizable with the same factor $(Z_{T}^{s})^{2}$, ensuring that $\kappa_{TX}^{s}=(\chi_{T}^{s}-\chi_{X}^{s})/(\chi_{T}^{s}+\chi_{X}^{s})$ is RG-invariant. Measuring this ratio in future studies will therefore provide a direct probe of full $U(1)_{A}$ effective restoration in the singlet sector. If the hierarchical picture is correct, $\kappa_{\omega_{T},h_{1T}}$ should remain non-zero well above $T_{c}^{ns}$ and approach zero only at a higher temperature $T_{c}^{s}\gg T_{c}^{ns}$.
6.3 Broader symmetry landscape and emergent phases above $T_{c}$
The hierarchical picture presented above aligns with growing evidence for multi-stage transitions above $T_{c}$. Beyond the familiar $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, recent lattice studies reveal emergent approximate symmetries absent in the classical QCD Lagrangian. In particular, approximate chiral-spin $SU(2)_{\mathrm{CS}}$ symmetry Glozman:2014mka observed in $N_{f}=2$ Rohrhofer:2019qwq ; Rohrhofer:2019qal , $N_{f}=2+1+1$ Chiu:2023hnm ; Chiu:2024jyz , and $N_{f}=2+1+1+1$ Chiu:2024bqx lattice QCD indicates that hadron-like states bound predominantly by chromoelectric interactions persist well above $T_{c}$.
It is important to distinguish this emergent symmetry from the usual chiral symmetries. For $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$, symmetry breaking decreases monotonically with temperature, and the symmetries themselves become effectively restored as partner degeneracy improves. These symmetries are always present in the Lagrangian — only their breaking weakens.
In contrast, $SU(2)_{\mathrm{CS}}$ chiral-spin symmetry is not present in the QCD Lagrangian. It emerges only in an intermediate window where chromoelectric binding dominates, typically from $T\sim 2T_{c}$ up to $T\sim 4T_{c}$ for light $(u,d)$ quarks. The window boundaries depend on two thresholds: $\epsilon_{\mathrm{CS}}$ for emergence at $T_{\mathrm{CS}}$, and $\epsilon_{\mathrm{CSF}}$ for fading at $T_{\mathrm{CSF}}>T_{\mathrm{CS}}$. Within this window, $SU(2)_{\mathrm{CS}}$ breaking decreases with $T$, allowing hadron-like states bound by chromoelectric interactions to exist. As temperature increases, thermal excitation energy eventually exceeds the chromoelectric binding energy, leading to quark deconfinement. This is manifested when one $SU(2)_{\mathrm{CS}}$ multiplet $(V_{k},T_{k},X_{k})$ merges with the $(P,S)$ multiplet of $U(1)_{A}$ symmetry, leaving only the usual chiral symmetries relevant. At this point, chiral-spin symmetry fades away — not simply becoming more broken, but physically disappearing due to thermal excitations.
Similarly, the observation of an infrared symmetric phase Alexandru:2019gdm around $T_{\mathrm{IR}}\sim 230$ MeV in $N_{f}=2+1$ lattice QCD Meng:2023nxf , which disappears by $T\gtrsim 300$ MeV, supports a correlated, quasi-hadronic medium in the window $T_{c}\lesssim T\lesssim 2T_{c}$. In this regime, hadrons become progressively delocalized, forming a dense fluid where both chromoelectric and chromomagnetic interactions remain active. The usual chiral multiplets are already restored in the nonsinglet sector, but chiral-spin symmetry is not yet manifest.
The gradual loss of binding across both regimes is consistent with the observed reduction in $U(1)_{A}$-breaking effects in connected channels (as measured by $\kappa_{PS}$ and $\kappa_{TX}$), while topological fluctuations—including center vortices tHooft:1977nqb and instanton-like objects—remain active until higher temperatures.
These emergent symmetry patterns are complemented by topological studies. Center-vortex analyses in $N_{f}=2+1$ QCD observe a pronounced change in vortex percolation at approximately $\gtrsim 2T_{c}$ Mickley:2024vkm . While vortex density drops substantially, a sparse network survives to higher temperatures, reinforcing that topological fluctuations — intimately connected to the $U(1)_{A}$ anomaly — remain non-negligible well into the deconfined regime. This provides a plausible mechanism for continued $U(1)_{A}$ breaking where nonsinglet chiral symmetry is largely restored. The smooth crossover implies no sharp “complete deconfinement” point; instead, the system evolves gradually from hadronic to quark–gluon plasma, with residual correlations possible even at $T\gg 2T_{c}$.
Taken together, these independent lines of evidence support the two-stage effective restoration scenario described in section 6.1. Stage 1 ($T\sim T_{c}^{ns}$) is characterized by effective restoration of nonsinglet symmetries, as demonstrated by the vanishing of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ in this work. Stage 2 ($T\sim T_{c}^{s}$) involves full effective restoration of singlet channels and suppression of topological fluctuations ($\chi_{t}\to 0$). Consistent comparison between lattice studies requires continuum extrapolation of each observable, as demonstrated here. Since the chiral transition at physical quark masses is a smooth crossover, any symmetry-breaking observable decreases gradually with temperature. The Stage 2 scenario remains to be conclusively established through future continuum-extrapolated measurements of tensor-singlet ratios and topological susceptibility.
6.4 Future computational improvements and extensions
6.4.1 Computational improvements and tensor singlet probes
To build upon this work, several computational improvements are planned:
  • Continuum extrapolation: We will perform simulations at several fixed temperatures, each with multiple lattice spacings to reduce systematic uncertainties in the continuum extrapolation of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$.
  • Finite-volume effects: We will employ lattices with larger spatial volumes ($40^{3}$ and $64^{3}$) to better distinguish between finite-size artifacts and genuine thermodynamic behavior.
  • Statistical precision: We will increase the number of gauge configurations in each ensemble to reduce statistical uncertainties in correlation function measurements.
  • Flavor-singlet tensor channels: Using stochastic all-to-all propagators with color–Dirac dilution, we will compute the singlet tensor ($\omega_{T}$) and axial-tensor ($h_{1T}$) susceptibilities. The tensor singlet ratio
    $\kappa_{\omega_{T},h_{1T}}\equiv\frac{\chi_{\omega_{T}}-\chi_{h_{1T}}}{\chi_{\omega_{T}}+\chi_{h_{1T}}},$
    where $\omega_{T,k}=\bar{q}\gamma_{4}\gamma_{k}q$ and $h_{1T,k}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$, provides a theoretically clean (power-divergence free) probe of $U(1)_{A}$ restoration in the singlet sector. This will enable a direct test of the two-stage $U(1)_{A}$ restoration scenario at $T_{c}^{s}\gg T_{c}^{ns}$.
  • Mixed singlet–nonsinglet channels: The mixed ratios $\kappa(\chi_{\omega_{T}},\chi_{b_{1}})$ and $\kappa(\chi_{h_{1T}},\chi_{\rho_{T}})$ probe $SU(2)_{L}\times SU(2)_{R}$ restoration between singlet and nonsinglet tensor operators, offering additional insight into the chiral symmetry realization pattern across singlet and nonsinglet flavor sectors.
6.4.2 Extension to flavor-singlet channels
Beyond the flavor-nonsinglet mesons studied here, the restoration patterns of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry and $U(1)_{A}$ axial symmetry can also be probed through flavor-singlet meson channels. To verify consistency in the continuum limit, we will evaluate all-to-all quark propagators using stochastic estimation. The stochastic sources will be constructed with $Z_{2}$ noise vectors, diluted in both color and Dirac space. Specifically, each noise vector will have support on a single color–Dirac component ($3\times 4$ in total), with entries randomly chosen as $\pm 1$. This dilution scheme reduces off-diagonal contamination and improves the signal-to-noise ratio. A practical challenge lies in controlling statistical uncertainties for disconnected diagrams. We will explore advanced variance-reduction techniques to achieve sufficient precision for the flavor-singlet channels, alongside direct measurements of $\chi_{t}$ using the overlap operator and index theorem to comprehensively map the symmetry restoration landscape.
6.5 RG interpretation and comparison with Pisarski–Wilczek analyses
The renormalization-group analysis of Pisarski-Wilczek (PW) Pisarski:1983ms shows that the order of the two-flavor chiral transition is controlled by the RG scaling of the $U(1)_{A}$-breaking operator at the $O(4)$ fixed point. If this operator is relevant, axial symmetry remains broken and the infrared flow is governed by the stable $O(4)$ Wilson–Fisher fixed point, allowing a second-order transition in the $O(4)$ universality class. If it is irrelevant, the effective symmetry enlarges to $U(2)_{L}\times U(2)_{R}$; the $\epsilon$-expansion of the corresponding Landau–Ginzburg theory reveals no infrared-stable fixed point, implying runaway RG flow and a fluctuation-induced first-order transition.
The PW scenario applies strictly to the chiral limit $(m_{q}=0)$ at the exact infrared critical fixed point and therefore characterizes universality classes, rather than specific observables at finite quark mass. In physical QCD with $m_{q}\neq 0$, where the transition is a smooth crossover and no true critical fixed point exists, axial symmetry restoration must be defined operationally. In this work, we define “effective restoration” through the vanishing of the $\kappa_{AB}$ ratio within statistical uncertainties (section 2.3). This condition is substantially weaker than the RG irrelevance criterion of PW: it does not imply enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ or a change in universality class, but only that symmetry-breaking effects are not resolved within the present observables and statistical precision.
With this distinction, our results reveal a two-stage pattern of operational symmetry restoration.
Stage 1 (nonsinglet sector): Near $T_{c}^{\mathrm{ns}}\sim 156$ MeV, the ratios $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ become consistent with zero, indicating operational effective restoration of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ in the nonsinglet sector. However, this temperature does not correspond to the critical temperature of the PW framework, which refers to the chiral limit and requires restoration of full chiral symmetry at the infrared fixed point. Since singlet channels remain nondegenerate, the universality-class arguments of PW do not directly apply at this stage.
Stage 2 (singlet sector): At a significantly higher temperature $T_{c}^{\mathrm{s}}\gg T_{c}^{\mathrm{ns}}$, the singlet ratio $\kappa_{\omega_{T},h_{1T}}$ and the topological susceptibility $\chi_{t}$ also become consistent with zero, extending operational effective restoration to both singlet and nonsinglet sectors. This defines a more complete restoration scale than the PW critical temperature, which is determined solely by chiral symmetry in the massless limit and does not require observable axial degeneracy. Nevertheless, even at $T_{c}^{\mathrm{s}}$, this operational restoration does not by itself establish enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ or determine the universality class, as explicit symmetry-breaking effects from finite quark masses and the anomaly remain present in the Lagrangian. Rather, $T_{c}^{\mathrm{s}}$ marks the temperature at which symmetry breaking becomes unresolvable in all measured channels, representing a stronger phenomenological restoration scale beyond the strict RG definition of criticality.
Pisarski and Wilczek (PW) Pisarski:1983ms identified two scenarios depending on whether the $U(1)_{A}$-breaking operator is relevant or irrelevant at the $O(4)$ fixed point, with the latter implying enlargement of the symmetry to $U(2)_{L}\times U(2)_{R}$ and a fluctuation-induced first-order transition. The subsequent renormalization-group analysis of Pelissetto and Vicari (PV) Pelissetto:2013hqa established that for $N_{f}=2$ this operator is marginally irrelevant at the $O(4)$ fixed point, implying approximate $O(4)$ scaling in the chiral limit near the PW critical temperature. Our results differ qualitatively from this scenario. At $T_{c}^{\mathrm{ns}}\sim 156$ MeV, we observe degeneracy in nonsinglet channels while singlet observables and the topological susceptibility remain nonvanishing, revealing a two-stage hierarchy of operational restoration. This structure does not follow from the fixed-point analysis, which applies strictly in the chiral limit and does not distinguish connected and disconnected contributions. Moreover, our temperatures $T_{c}^{\mathrm{ns}}$ and $T_{c}^{\mathrm{s}}$ are crossover scales in physical QCD rather than critical points, and the persistence of $\chi_{t}>0$ reflects the continued contribution of disconnected diagrams and the mass-dependent amplification of axial-anomaly effects. Thus, while marginal irrelevance provides an important conceptual backdrop, the hierarchical restoration observed here represents a distinct finite-mass phenomenon beyond the scope of the chiral-limit RG framework.
This interpretation is supported by complementary lattice studies probing chiral symmetry restoration from different perspectives. In the $N_{f}=2$ chiral limit, continuum-controlled simulations using a many-flavor approach find that the apparent first-order behavior on coarse lattices is a discretization artifact, with the continuum transition consistent with second order and compatible with $O(N)$ scaling Klinger:2025xxb . At the physical point in $N_{f}=2+1+1$ QCD, independent scaling analysis with twisted-mass Wilson fermions Kotov:2021rah shows quantitative agreement with 3D $O(4)$ scaling over a broad temperature range, indicating that approximate critical scaling emerges already at nonzero quark mass.
Within this framework, our $\kappa_{AB}$ observables provide the first continuum-extrapolated, quantitative evidence for Stage 1, namely the operational restoration of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ in the nonsinglet (connected) sector near $T_{c}$. It is worth noting that other ratios such as $\kappa_{\sigma,\pi}$ and $\kappa_{\eta,\delta}$, which involve singlet channels, are also formally RG-invariant, as the underlying renormalization constants satisfy $Z_{S}^{ns}=Z_{P}^{ns}=Z_{S}^{s}=Z_{P}^{s}$ (see section 2.4). However, in practice, even after contact term and vacuum subtractions, the singlet susceptibilities $\chi_{\sigma}$ and $\chi_{\eta}$ retain finite mixing with the topological susceptibility $\chi_{t}$, which must be disentangled to obtain fully renormalized quantities. This additional complication, together with the noise from disconnected diagrams, makes these ratios less accessible in the present study.
Together, these results support a coherent picture in which the axial anomaly becomes progressively less visible near the chiral crossover. The chiral-limit studies establish the relevance of $O(N)$-type critical behavior at the true fixed point, while physical-mass simulations demonstrate that approximate $O(4)$-like scaling and chiral partner degeneracy appear as precursor phenomena. Our observation of nonsinglet degeneracy near $T_{c}^{\mathrm{ns}}$ is consistent with this framework, while the delayed restoration in singlet channels reveals additional finite-mass and anomaly-driven effects beyond the strict chiral-limit universality-class description.
Functional renormalization-group analyses of $(2+1)$-flavor QCD Braun:2020ada similarly find that axial-anomaly effects persist throughout a broad crossover region. While nonsinglet observables exhibit degeneracy patterns consistent with approximate $O(4)$-like symmetry, singlet channels remain sensitive to anomaly-driven disconnected contributions. As a result, the fully symmetric $U(2)_{L}\times U(2)_{R}$ scenario—requiring both $m_{q}=0$ and exact irrelevance of axial breaking at the infrared fixed point—is not realized at the physical crossover temperature and is therefore not directly applicable to our simulation. Instead, the system follows a hierarchical restoration pattern consistent with our two-stage picture.
Establishing Stage 2, corresponding to full operational restoration including singlet channels, requires both $\kappa_{\omega_{T},h_{1T}}\to 0$ and $\chi_{t}\to 0$ within uncertainties. Existing lattice studies show that topological fluctuations persist up to $T\gtrsim 2T_{c}$ (see e.g., Petreczky:2016vrs ; Borsanyi:2016ksw ; Chen:2022fid ; Athenodorou:2022aay ; Kotov:2025ilm ), indicating that anomaly effects remain active well above the crossover. However, a definitive continuum-limit determination combining chirally symmetric fermions and overlap-based topology measurements for both sea and valence sectors is still lacking. Such calculations are necessary to determine whether and at what temperature full effective restoration occurs.
Importantly, our results do not contradict the PW analysis, but instead probe a different regime. The PW framework applies strictly to the chiral limit at the infrared critical fixed point, where universal $O(4)$ scaling would govern all channels. By contrast, our study concerns physical QCD with nonzero quark masses, where symmetry restoration occurs through a crossover rather than a true critical transition. The observed degeneracy of nonsinglet chiral partners near $T_{c}$ is a necessary precursor for $O(4)$-like scaling, but does not by itself establish the associated universality class. The behavior of the singlet sector—and whether its eventual operational restoration at $T_{c}^{\mathrm{s}}$ bears any connection to the $U(2)_{L}\times U(2)_{R}$ scenario considered in the chiral-limit analysis—remains an open question. Future continuum-extrapolated measurements of $\kappa_{TX}^{s}$ will directly probe the anomaly’s role in the singlet sector and further constrain the underlying symmetry-restoration mechanism.
6.6 Concluding remarks
The symmetry-strength parameter $\kappa_{AB}$ introduced here provides a flexible and systematically improvable framework for quantifying symmetry breaking in QCD. Applied to finite-temperature QCD, it leads to several central conclusions:
1. RG-invariant diagnostics: The $\kappa_{AB}$ ratios furnish renormalization-group invariant, scheme-independent measures of symmetry breaking across different operator channels, allowing direct comparison between lattice spacings and formulations.
2. Continuum convergence: Although finite-$a$ effects produce channel-dependent hierarchies, all nonsinglet $\kappa_{AB}$ values converge consistently in the continuum limit, indicating a common restoration scale for different manifestations of chiral and axial symmetry in the connected sector.
3. Hierarchical restoration: Both $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ exhibit a two-stage restoration pattern in the crossover regime: first in the nonsinglet sector around $T_{c}^{\mathrm{ns}}\sim 156$ MeV, and only at significantly higher temperature in the singlet sector, where suppression of topological fluctuations becomes essential.
[…]
Natural extensions include simulations at lighter quark masses to probe scaling behavior more directly, applications to additional operator and flavor sectors, and systematic comparison with functional renormalization-group and effective-model studies. More broadly, the $\kappa_{AB}$ program establishes a quantitative benchmark for assessing symmetry realization in QCD and clarifies the relation between lattice observables and continuum effective descriptions of the chiral transition.
Combined with future continuum-extrapolated determinations of tensor singlet channels and overlap-fermion-based measurements of $\chi_{t}$, this framework offers a comprehensive strategy for mapping the complete restoration pattern of chiral and axial symmetries in QCD.
Appendix A Meson Operator Notation
Table 5: Summary of meson operators and their corresponding regularized susceptibilities in $N_{f}=2$ QCD. The notation follows eqs. (7)–(15) in the main text. Singlet ($s$) and nonsinglet (nsns) channels are distinguished, with the latter involving only connected diagrams. Tensor operators $T_{k}$ and $X_{k}$ ($k=1,2,3$) are defined with spatial index $k$.
Meson Operator Channel Flavor Susceptibility
$\pi^{a}$ $\bar{q}\gamma_{5}t^{a}q$ Pseudoscalar ($P$) Nonsinglet ($ns$) $\chi_{\pi}=\chi_{P}^{ns}$
$\delta^{a}$ $\bar{q}t^{a}q$ Scalar ($S$) Nonsinglet ($ns$) $\chi_{\delta}=\chi_{S}^{ns}$
$\sigma$ $\bar{q}q$ Scalar ($S$) Singlet ($s$) $\chi_{\sigma}=\chi_{S}^{s}$
$\eta$ ($N_{f}=2$) $\bar{q}\gamma_{5}q$ Pseudoscalar ($P$) Singlet ($s$) $\chi_{\eta}=\chi_{P}^{s}$
$\rho^{a}$ $\bar{q}\gamma_{i}t^{a}q$ Vector ($V_{i}$) Nonsinglet ($ns$) $\chi_{\rho}=\chi_{V_{i}}^{ns}$
$a_{1}^{a}$ $\bar{q}\gamma_{5}\gamma_{i}t^{a}q$ Axial-vector ($A_{i}$) Nonsinglet ($ns$) $\chi_{a_{1}}=\chi_{A_{i}}^{ns}$
$\omega$ $\bar{q}\gamma_{i}q$ Vector ($V_{i}$) Singlet ($s$) $\chi_{\omega}=\chi_{V_{i}}^{s}$
$h_{1}$ $\bar{q}\gamma_{5}\gamma_{i}q$ Axial-vector ($A_{i}$) Singlet ($s$) $\chi_{h_{1}}=\chi_{A_{i}}^{s}$
$\rho_{T}^{a}$ $\bar{q}\gamma_{4}\gamma_{i}t^{a}q$ Tensor ($T_{i}$) Nonsinglet ($ns$) $\chi_{\rho_{T}}=\chi_{T_{i}}^{ns}$
$b_{1}^{a}$ $\bar{q}\gamma_{5}\gamma_{4}\gamma_{i}t^{a}q$ Axial-tensor ($X_{i}$) Nonsinglet ($ns$) $\chi_{b_{1}}=\chi_{X_{i}}^{ns}$
$\omega_{T}$ $\bar{q}\gamma_{4}\gamma_{i}q$ Tensor ($T_{i}$) Singlet ($s$) $\chi_{\omega_{T}}=\chi_{T_{i}}^{s}$
$h_{1T}$ $\bar{q}\gamma_{5}\gamma_{4}\gamma_{i}q$ Axial-tensor ($X_{i}$) Singlet ($s$) $\chi_{h_{1T}}=\chi_{X_{i}}^{s}$
  • For $N_{f}=2$ flavor, the generators are $t^{0}=\mbox{1\hskip-3.41432ptI}/2$ (for flavor singlet) and $t^{a}=\tau^{a}/2$ ($a=1,2,3$ for flavor nonsinglets), where 1I is the $2\times 2$ identity matrix and $\tau^{a}$ are the Pauli matrices. They satisfy the normalization condition ${\rm tr}(t^{a}t^{b})=\delta^{ab}/2$ for all $a,b=0,1,2,3$.
  • The flavor‑singlet generator $t^{0}$ is always suppressed in our notation for singlet operators; e.g., we write $\bar{q}q$ to mean $\bar{q}t^{0}q$, and similarly $\bar{q}\gamma_{5}q$ for $\bar{q}\gamma_{5}t^{0}q$.
  • Spatial index $i=1,2,3$ (or $k=1,2,3$) is implicit for vector and tensor channels.
  • Singlet operators ($\sigma$, $\eta$, $\omega$, $h_{1}$, $\omega_{T}$, $h_{1T}$) involve disconnected diagrams and require careful subtraction of power divergences.
  • Nonsinglet operators involve only connected diagrams and have cleaner renormalization properties.
  • The tensor representations $\rho_{T}^{a}$ (via $T_{i}$) and $b_{1}^{a}$ (via $X_{i}$) probe $U(1)_{A}$ through a different Dirac structure than the scalar–pseudoscalar pair ($\delta$$\pi$).
  • The vector singlet $\omega$ and axial-vector singlet $h_{1}$ are the $I=0$ partners under $U(1)_{A}$, while the tensor singlet $\omega_{T}$ and axial-tensor singlet $h_{1T}$ play analogous roles in the tensor channel.
Appendix B Tables of results
In this appendix, we provide numerical data of the figures in the main text. Specifically, the data for the time-correlation functions, regularized susceptibilities, and RG-invariant symmetry ratios for $u,d$ off-diagonal flavor-nonsinglet mesons are presented in tabular form, corresponding to figures 1, 2, and 3, respectively.
The numerical values for the time-correlation functions $C_{\Gamma}(t)$ of the $\bar{u}\Gamma d$ on the $32^{3}\times 16$ lattice, shown in figure 1, are tabulated in tables 68 for the $(P,S,V,A,T,X)$ channels, for three lattice spacings $a$ = (0.075, 0.069, 0.064) fm.
Numerical results for the regularized susceptibilities $(\chi_{S},\chi_{P},\chi_{V},\chi_{A},\chi_{T},\chi_{X})$, presented in figure 2, are listed in tables 911. These tables also include the corresponding RG-invariant symmetry ratios $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$, shown in figures 3. Data are provided for the same three lattice spacings across twelve temperatures ranging from 164 to 385 MeV.
Statistical uncertainties are estimated using the jackknife method with a bin size of 10101515 configurations of which the statistical error saturates. The numerical values are quoted with more significant figures than strictly warranted by the precision. This choice is intentional, to clearly resolve subtle differences—particularly between symmetry partners such as $\chi_{V}$ and $\chi_{A}$ —that might otherwise be obscured within the quoted error bars.
$t/a$ $C_{P}(t)$ $C_{S}(t)$ $C_{V}(t)$ $C_{A}(t)$ $C_{T}(t)$ $C_{X}(t)$
1 $3.926(15)\times 10^{-2}$ $3.919(15)\times 10^{-2}$ $2.9286(88)\times 10^{-2}$ $2.9285(88)\times 10^{-2}$ $2.1937(41)\times 10^{-2}$ $2.1937(41)\times 10^{-2}$
2 $6.462(91)\times 10^{-3}$ $6.397(93)\times 10^{-3}$ $3.291(23)\times 10^{-3}$ $3.291(23)\times 10^{-3}$ $1.971(14)\times 10^{-3}$ $1.971(14)\times 10^{-3}$
3 $2.611(85)\times 10^{-3}$ $2.551(84)\times 10^{-3}$ $8.317(75)\times 10^{-4}$ $8.314(75)\times 10^{-4}$ $4.387(50)\times 10^{-4}$ $4.387(50)\times 10^{-4}$
4 $1.620(84)\times 10^{-3}$ $1.563(83)\times 10^{-3}$ $3.299(35)\times 10^{-4}$ $3.297(35)\times 10^{-4}$ $1.682(26)\times 10^{-4}$ $1.682(26)\times 10^{-4}$
5 $238(83)\times 10^{-3} 1.1$$184(82)\times 10^{-3}$ $1.754(23)\times 10^{-4}$ $1.752(23)\times 10^{-4}$ $9.16(19)\times 10^{-5}$ $9.16(19)\times 10^{-5}$
6 $1.062(82)\times 10^{-3}$ $012(81)\times 10^{-3} 1.1$$60(17)\times 10^{-4} 1.1$$159(17)\times 10^{-4}$ $6.35(16)\times 10^{-5}$ $6.35(16)\times 10^{-5}$
7 $9.82(81)\times 10^{-4}$ $9.34(81)\times 10^{-4}$ $9.19(16)\times 10^{-5}$ $9.18(16)\times 10^{-5}$ $5.24(14)\times 10^{-5}$ $5.25(15)\times 10^{-5}$
8 $9.61(80)\times 10^{-4}$ $9.13(81)\times 10^{-4}$ $8.52(16)\times 10^{-5}$ $8.51(16)\times 10^{-5}$ $4.92(14)\times 10^{-5}$ $4.93(14)\times 10^{-5}$
9 $9.90(79)\times 10^{-4}$ $9.41(80)\times 10^{-4}$ $9.24(17)\times 10^{-5}$ $9.23(17)\times 10^{-5}$ $5.23(17)\times 10^{-5}$ $5.24(18)\times 10^{-5}$
1 $1.077(77)\times 10^{-3}$ $026(79)\times 10^{-3} 1.1$$76(18)\times 10^{-4} 1.1$$175(18)\times 10^{-4}$ $6.35(19)\times 10^{-5}$ $6.35(19)\times 10^{-5}$
11 $1.259(75)\times 10^{-3}$ $1.204(78)\times 10^{-3}$ $1.775(21)\times 10^{-4}$ $1.774(21)\times 10^{-4}$ $9.14(19)\times 10^{-5}$ $9.13(19)\times 10^{-5}$
12 $1.645(72)\times 10^{-3}$ $1.585(75)\times 10^{-3}$ $326(32)\times 10^{-4} 3.3$$324(32)\times 10^{-4}$ $1.681(25)\times 10^{-4}$ $1.681(25)\times 10^{-4}$
13 $2.645(71)\times 10^{-3}$ $2.579(75)\times 10^{-3}$ $8.384(53)\times 10^{-4}$ $8.381(53)\times 10^{-4}$ $425(42)\times 10^{-4} 4.4$$424(42)\times 10^{-4}$
14 $6.521(74)\times 10^{-3}$ $6.450(80)\times 10^{-3}$ $322(17)\times 10^{-3} 3.3$$322(17)\times 10^{-3}$ $1.996(10)\times 10^{-3}$ $1.996(11)\times 10^{-3}$
15 $3.940(12)\times 10^{-2}$ $3.933(11)\times 10^{-2}$ $2.9396(71)\times 10^{-2}$ $9395(71)\times 10^{-2} 2.2$$020(44)\times 10^{-2} 2.2$$2019(44)\times 10^{-2}$
Table 6: Time-correlation function $C_{\Gamma}(t)$ of $\bar{u}\Gamma d$ for $P,S,V,A,T,X$ channels on the $32^{3}\times 16$ lattice with lattice spacing aa=0.075 fm.
$t/a$ $C_{P}(t)$ $C_{S}(t)$ $C_{V}(t)$ $C_{A}(t)$ $C_{T}(t)$ $C_{X}(t)$
1 $3.944(11)\times 10^{-2}$ $3.943(11)\times 10^{-2}$ $2.9601(71)\times 10^{-2}$ $9601(71)\times 10^{-2} 2.2$$085(44)\times 10^{-2} 2.2$$2085(44)\times 10^{-2}$
2 $6.304(64)\times 10^{-3}$ $6.295(63)\times 10^{-3}$ $377(15)\times 10^{-3} 3.3$$376(15)\times 10^{-3}$ $2.0059(93)\times 10^{-3}$ $2.0057(93)\times 10^{-3}$
3 $2.368(59)\times 10^{-3}$ $2.360(58)\times 10^{-3}$ $8.639(57)\times 10^{-4}$ $8.638(57)\times 10^{-4}$ $479(37)\times 10^{-4} 4.4$$478(37)\times 10^{-4}$
4 $1.349(55)\times 10^{-3}$ $1.342(54)\times 10^{-3}$ $3.463(26)\times 10^{-4}$ $3.462(26)\times 10^{-4}$ $1.719(19)\times 10^{-4}$ $1.718(19)\times 10^{-4}$
5 $9.60(52)\times 10^{-4}$ $9.54(51)\times 10^{-4}$ $1.850(16)\times 10^{-4}$ $1.849(16)\times 10^{-4}$ $9.28(14)\times 10^{-5}$ $9.27(14)\times 10^{-5}$
6 $83(50)\times 10^{-4} 7.7$$77(49)\times 10^{-4}$ $22(13)\times 10^{-4} 1.221$$21(13)\times 10^{-4}$ $6.34(12)\times 10^{-5}$ $6.33(12)\times 10^{-5}$
7 $7.01(50)\times 10^{-4}$ $6.95(49)\times 10^{-4}$ $9.63(12)\times 10^{-5}$ $9.62(12)\times 10^{-5}$ $5.14(16)\times 10^{-5}$ $5.14(16)\times 10^{-5}$
8 $6.77(50)\times 10^{-4}$ $6.70(49)\times 10^{-4}$ $8.93(15)\times 10^{-5}$ $8.92(15)\times 10^{-5}$ $4.81(16)\times 10^{-5}$ $4.80(16)\times 10^{-5}$
9 $7.02(51)\times 10^{-4}$ $6.95(50)\times 10^{-4}$ $9.71(19)\times 10^{-5}$ $9.70(19)\times 10^{-5}$ $5.16(17)\times 10^{-5}$ $5.15(17)\times 10^{-5}$
10 $87(52)\times 10^{-4} 7.7$$79(51)\times 10^{-4}$ $1.239(23)\times 10^{-4}$ $1.239(23)\times 10^{-4}$ $6.40(17)\times 10^{-5}$ $6.39(17)\times 10^{-5}$
11 $9.71(54)\times 10^{-4}$ $9.61(53)\times 10^{-4}$ $1.877(31)\times 10^{-4}$ $1.876(31)\times 10^{-4}$ $9.43(14)\times 10^{-5}$ $9.42(14)\times 10^{-5}$
12 $1.363(58)\times 10^{-3}$ $1.352(57)\times 10^{-3}$ $3.489(41)\times 10^{-4}$ $3.488(41)\times 10^{-4}$ $1.738(22)\times 10^{-4}$ $1.738(22)\times 10^{-4}$
13 $2.377(65)\times 10^{-3}$ $2.365(63)\times 10^{-3}$ $8.631(68)\times 10^{-4}$ $8.630(68)\times 10^{-4}$ $495(40)\times 10^{-4} 4.4$$494(40)\times 10^{-4}$
14 $6.28(11)\times 10^{-3}$ $6.26(10)\times 10^{-3}$ $356(19)\times 10^{-3} 3.3$$356(19)\times 10^{-3}$ $1.998(11)\times 10^{-3}$ $1.998(11)\times 10^{-3}$
15 $3.927(11)\times 10^{-2}$ $3.926(11)\times 10^{-2}$ $2.9478(74)\times 10^{-2}$ $9478(74)\times 10^{-2} 2.2$$017(45)\times 10^{-2} 2.2$$2017(45)\times 10^{-2}$
Table 7: Time-correlation function $C_{\Gamma}(t)$ of $\bar{u}\Gamma d$ for $P,S,V,A,T,X$ channels on the $32^{3}\times 16$ lattice with lattice spacing aa=0.069 fm.
$t/a$ $C_{P}(t)$ $C_{S}(t)$ $C_{V}(t)$ $C_{A}(t)$ $C_{T}(t)$ $C_{X}(t)$
1 $3.9219(97)\times 10^{-2}$ $3.9215(97)\times 10^{-2}$ $2.9480(64)\times 10^{-2}$ $2.9480(64)\times 10^{-2}$ $2.1966(39)\times 10^{-2}$ $2.1966(39)\times 10^{-2}$
2 $6.238(31)\times 10^{-3}$ $6.234(31)\times 10^{-3}$ $389(15)\times 10^{-3} 3.3$$389(15)\times 10^{-3}$ $2.0051(86)\times 10^{-3}$ $2.0051(86)\times 10^{-3}$
3 $2.307(32)\times 10^{-3}$ $2.304(32)\times 10^{-3}$ $8.733(45)\times 10^{-4}$ $8.732(45)\times 10^{-4}$ $4.506(28)\times 10^{-4}$ $4.506(28)\times 10^{-4}$
4 $1.287(27)\times 10^{-3}$ $1.284(27)\times 10^{-3}$ $3.508(21)\times 10^{-4}$ $3.508(21)\times 10^{-4}$ $1.735(16)\times 10^{-4}$ $1.735(16)\times 10^{-4}$
5 $8.96(26)\times 10^{-4}$ $8.93(25)\times 10^{-4}$ $1.861(13)\times 10^{-4}$ $1.861(13)\times 10^{-4}$ $9.34(17)\times 10^{-5}$ $9.34(17)\times 10^{-5}$
6 $7.17(25)\times 10^{-4}$ $7.14(25)\times 10^{-4}$ $1.210(11)\times 10^{-4}$ $1.210(11)\times 10^{-4}$ $6.34(17)\times 10^{-5}$ $6.34(17)\times 10^{-5}$
7 $6.37(25)\times 10^{-4}$ $6.34(25)\times 10^{-4}$ $9.36(12)\times 10^{-5}$ $9.36(12)\times 10^{-5}$ $5.11(16)\times 10^{-5}$ $5.10(16)\times 10^{-5}$
8 $6.16(26)\times 10^{-4}$ $6.13(26)\times 10^{-4}$ $8.56(12)\times 10^{-5}$ $8.55(12)\times 10^{-5}$ $4.74(21)\times 10^{-5}$ $4.74(21)\times 10^{-5}$
9 $6.45(27)\times 10^{-4}$ $6.42(27)\times 10^{-4}$ $9.24(16)\times 10^{-5}$ $9.24(16)\times 10^{-5}$ $5.04(21)\times 10^{-5}$ $5.04(21)\times 10^{-5}$
10 $7.32(28)\times 10^{-4}$ $7.29(28)\times 10^{-4}$ $186(16)\times 10^{-4} 1.1$$186(17)\times 10^{-4}$ $6.22(22)\times 10^{-5}$ $6.22(22)\times 10^{-5}$
11 $9.15(29)\times 10^{-4}$ $9.12(28)\times 10^{-4}$ $1.825(19)\times 10^{-4}$ $1.825(19)\times 10^{-4}$ $9.19(25)\times 10^{-5}$ $9.19(25)\times 10^{-5}$
12 $1.307(29)\times 10^{-3}$ $1.304(29)\times 10^{-3}$ $3.465(22)\times 10^{-4}$ $3.465(22)\times 10^{-4}$ $1.720(24)\times 10^{-4}$ $1.720(24)\times 10^{-4}$
13 $2.323(30)\times 10^{-3}$ $2.319(29)\times 10^{-3}$ $8.692(46)\times 10^{-4}$ $8.692(46)\times 10^{-4}$ $4.508(35)\times 10^{-4}$ $4.507(35)\times 10^{-4}$
14 $6.256(29)\times 10^{-3}$ $6.253(29)\times 10^{-3}$ $396(13)\times 10^{-3} 3.3$$395(13)\times 10^{-3}$ $2.0141(80)\times 10^{-3}$ $2.0141(80)\times 10^{-3}$
15 $3.9292(82)\times 10^{-2}$ $3.9288(82)\times 10^{-2}$ $2.9539(54)\times 10^{-2}$ $9538(54)\times 10^{-2} 2.2$$012(33)\times 10^{-2} 2.2$$2012(33)\times 10^{-2}$
Table 8: Time-correlation function $C_{\Gamma}(t)$ of $\bar{u}\Gamma d$ for $P,S,V,A,T,X$ channels on the $32^{3}\times 16$ lattice with lattice spacing aa=0.064 fm.
$T[\text{MeV}]$ $a[\text{fm}]$ $\chi_{S}$ $\chi_{P}$ $\kappa_{PS}$
164 0.075 5.34324e-2 $\pm$ 6.014e-4 5.38707e-2 $\pm$ 5.882e-4 4.0846e-3 $\pm$ 1.0146e-3
179 0.069 5.20737e-2 $\pm$ 3.699e-4 5.21386e-2 $\pm$ 3.797e-4 6.2278e-4 $\pm$ 2.9024e-4
192 0.064 5.15927e-2 $\pm$ 1.981e-4 5.16162e-2 $\pm$ 1.985e-4 2.2824e-4 $\pm$ 6.641e-5
219 0.075 5.02095e-2 $\pm$ 1.606e-4 5.02886e-2 $\pm$ 1.566e-4 7.8791e-4 $\pm$ 1.8618e-4
238 0.069 5.01252e-2 $\pm$ 2.235e-4 5.01403e-2 $\pm$ 2.260e-4 1.5054e-4 $\pm$ 3.545e-5
257 0.064 4.96265e-2 $\pm$ 1.391e-4 4.96309e-2 $\pm$ 1.389e-4 4.4259e-5 $\pm$ 2.0238e-5
263 0.075 4.79896e-2 $\pm$ 1.793e-4 4.80188e-2 $\pm$ 1.790e-4 3.0421e-4 $\pm$ 5.490e-5
286 0.069 4.80414e-2 $\pm$ 1.190e-4 4.80473e-2 $\pm$ 1.193e-4 6.1439e-5 $\pm$ 1.1913e-5
308 0.064 4.81840e-2 $\pm$ 1.089e-4 4.81862e-2 $\pm$ 1.089e-4 2.3046e-5 $\pm$ 5.357e-6
328 0.075 4.69035e-2 $\pm$ 1.273e-4 4.69140e-2 $\pm$ 1.259e-4 1.1254e-4 $\pm$ 4.477e-5
357 0.069 4.68727e-2 $\pm$ 1.649e-4 4.68751e-2 $\pm$ 1.657e-4 2.5877e-5 $\pm$ 6.983e-6
385 0.064 4.69490e-2 $\pm$ 1.166e-4 4.69499e-2 $\pm$ 1.166e-4 9.5128e-6 $\pm$ 3.3590e-6
Table 9: Scalar and pseudoscalar susceptibilities $\chi_{S}$, $\chi_{P}$ and the RG-invariant symmetry ratio $\kappa_{PS}$ for three lattice spacings and twelve temperatures.
$T[\text{MeV}]$ $a[\text{fm}]$ $\chi_{V}$ $\chi_{A}$ $\kappa_{VA}$
164 0.075 3.416665e-2 $\pm$ 8.403e-5 3.416499e-2 $\pm$ 8.402e-5 2.4248e-5 $\pm$ 2.080e-6
179 0.069 3.456805e-2 $\pm$ 7.824e-5 3.456716e-2 $\pm$ 7.824e-5 1.2819e-5 $\pm$ 1.806e-6
192 0.064 3.456541e-2 $\pm$ 6.396e-5 3.456498e-2 $\pm$ 6.395e-5 6.2764e-6 $\pm$ 4.664e-7
219 0.075 3.414924e-2 $\pm$ 9.063e-5 3.414868e-2 $\pm$ 9.063e-5 8.1562e-6 $\pm$ 9.217e-7
238 0.069 3.448263e-2 $\pm$ 6.622e-5 3.448232e-2 $\pm$ 6.622e-5 4.5307e-6 $\pm$ 2.865e-7
257 0.064 3.460791e-2 $\pm$ 6.813e-5 3.460776e-2 $\pm$ 6.813e-5 2.2422e-6 $\pm$ 3.596e-7
263 0.075 3.412941e-2 $\pm$ 8.999e-5 3.412913e-2 $\pm$ 8.999e-5 4.0453e-6 $\pm$ 3.488e-7
286 0.069 3.446740e-2 $\pm$ 5.640e-5 3.446724e-2 $\pm$ 5.640e-5 2.2984e-6 $\pm$ 8.424e-8
308 0.064 3.472431e-2 $\pm$ 7.547e-5 3.472421e-2 $\pm$ 7.547e-5 1.4126e-6 $\pm$ 5.159e-8
328 0.075 3.384798e-2 $\pm$ 7.409e-5 3.384784e-2 $\pm$ 7.409e-5 2.1217e-6 $\pm$ 2.826e-7
357 0.069 3.449452e-2 $\pm$ 8.048e-5 3.449444e-2 $\pm$ 8.048e-5 1.2518e-6 $\pm$ 1.138e-7
385 0.064 3.478435e-2 $\pm$ 7.885e-5 3.478430e-2 $\pm$ 7.885e-5 6.8488e-7 $\pm$ 1.704e-8
Table 10: Vector and axial-vector susceptibilities $\chi_{V}$, $\chi_{A}$ and the RG-invariant symmetry ratio $\kappa_{VA}$ for three lattice spacings and twelve temperatures.
$T[\text{MeV}]$ $a[\text{fm}]$ $\chi_{T}$ $\chi_{X}$ $\kappa_{TX}$
164 0.075 2.480619e-2 $\pm$ 6.076e-5 2.480472e-2 $\pm$ 6.071e-5 2.9511e-5 $\pm$ 1.3839e-5
179 0.069 2.489018e-2 $\pm$ 4.489e-5 2.488948e-2 $\pm$ 4.490e-5 1.3890e-5 $\pm$ 3.475e-6
192 0.064 2.487534e-2 $\pm$ 3.871e-5 2.487512e-2 $\pm$ 3.871e-5 4.4736e-6 $\pm$ 5.302e-7
219 0.075 2.490372e-2 $\pm$ 5.385e-5 2.490318e-2 $\pm$ 5.379e-5 1.0886e-5 $\pm$ 3.168e-6
238 0.069 2.498616e-2 $\pm$ 4.239e-5 2.498592e-2 $\pm$ 4.239e-5 4.9238e-6 $\pm$ 3.050e-7
257 0.064 2.508379e-2 $\pm$ 3.896e-5 2.508370e-2 $\pm$ 3.896e-5 1.9293e-6 $\pm$ 1.303e-7
263 0.075 2.516102e-2 $\pm$ 4.261e-5 2.516079e-2 $\pm$ 4.261e-5 4.4966e-6 $\pm$ 9.353e-7
286 0.069 2.526107e-2 $\pm$ 3.355e-5 2.526093e-2 $\pm$ 3.355e-5 2.7641e-6 $\pm$ 2.086e-7
308 0.064 2.536585e-2 $\pm$ 4.611e-5 2.536578e-2 $\pm$ 4.611e-5 1.2422e-6 $\pm$ 1.613e-7
328 0.075 2.537214e-2 $\pm$ 4.411e-5 2.537201e-2 $\pm$ 4.411e-5 2.4743e-6 $\pm$ 6.449e-7
357 0.069 2.561389e-2 $\pm$ 5.092e-5 2.561378e-2 $\pm$ 5.092e-5 2.0030e-6 $\pm$ 4.424e-7
385 0.064 2.578521e-2 $\pm$ 4.603e-5 2.578517e-2 $\pm$ 4.603e-5 7.8727e-7 $\pm$ 2.746e-8
Table 11: Tensor-vector and axial-tensor-vector susceptibilities $\chi_{T}$, $\chi_{X}$ and the RG-invariant symmetry ratio $\kappa_{TX}$ for three lattice spacings and twelve temperatures.
Acknowledgements.
We are grateful to Academia Sinica Grid Computing Center and National Center for High Performance Computing for the computer time and facilities. This work was supported by the National Science and Technology Council (Grants No. 108-2112-M-003-005, No. 109-2112-M-003-006, No. 110-2112-M-003-009), and Academia Sinica Grid Computing Centre (Grant No. AS-CFII-112-103).